### v2

# Instant Runoff Voting and the Reinforcement Paradox

Abstract.
We analyze the susceptibility of instant runoff voting (IRV) to a lesser-studied paradox known as a reinforcement paradox, which occurs when candidate $X$ wins under IRV in two distinct elections but $X$ loses in the combined election formed by merging the ballots from the two elections. For three-candidate IRV elections we provide necessary and sufficient conditions under which there exists a partition of the ballot set into two sets of ballots such that a given losing candidate wins each of the sub-elections. Using these conditions, we use Monte Carlo simulations to estimate the frequency with which such partitions exist under various models of voter behavior. We also analyze the frequency with which the paradox occurs in a large dataset of real-world ranked-choice elections to provide empirical probabilities. Our general finding is that IRV is highly susceptible to this paradox in three-candidate elections.
Key words and phrases:
instant runoff voting, ranked choice voting, reinforcement paradox, consistency, empirical results
1. Introduction
The voting method of instant runoff voting (IRV), also referred to as the alternative vote, the plurality elimination rule, and colloquially as “ranked-choice voting” in the United States, is famously susceptible to many classical voting paradoxes [14, 32, 41]. In this article we analyze the susceptibility of IRV to a lesser-studied paradox known as a reinforcement paradox, which occurs when candidate $X$ wins under IRV in two distinct elections but $X$ loses in the combined election formed by merging the ballots from the two elections. For three-candidate IRV elections we provide necessary and sufficient conditions under which there exists a partition of the ballot set into two subsets of ballots such that a given losing candidate wins each of the sub-elections. Using these conditions, we use Monte Carlo simulations to estimate the frequency with which such partitions exist under various models of voter behavior. We then analyze the frequency of the paradox in a large dataset of real-world ranked-choice elections to provide empirical probabilities. By characterizing the three-candidate IRV elections which demonstrate the reinforcement paradox and estimating its frequency, we view our work as a contribution to the larger agenda laid out by Felsenthal in [12].
Reinforcement paradoxes have received attention in the social choice literature due to their interesting and surprising nature, much like other paradoxes to which IRV is susceptible, such as upward or downward monotonicity paradoxes. However, unlike monotonicity paradoxes, IRV’s susceptibility to reinforcement paradoxes has direct administrative implications: all ballots must be transmitted—either electronically or physically—to a central counting location. In contrast, a simpler system like plurality allows each precinct to report vote totals individually, and a winner can be determined by summing these across precincts. Due to IRV’s vulnerability to reinforcement paradoxes, precinct-by-precinct counting is not possible. As a result, IRV ballot tabulation may experience administrative delays that would not occur with methods like plurality. Thus, a study of reinforcement paradoxes is motivated by practical as well as theoretical considerations.
Our work in this paper connects to three previous strands of the literature concerning IRV and voting paradoxes. The first gives conditions on a preference profile which characterize when an election demonstrates a given paradox. Examples are Lepelley et al. [21], Miller [29], and Ornstein and Norman [33], which characterize the three-candidate profiles which demonstrate an upward or downward monotonicity paradox under IRV. Similarly, Graham-Squire [16] characterizes three-candidate profiles which demonstrate a no-show paradox under IRV, including profiles which contain partial ballots. Kamwa et al. [18] study a variety of voting rules, but in particular provide conditions under which three-candidate profiles produce paradoxes such as negative participation paradoxes under IRV.
The second strand of literature typically uses the conditions given by the works cited above to estimate the frequency with which a given paradox occurs under IRV. For example, Ornstein and Norman [33] estimate the frequency of monotonicity failures in IRV elections when the ballot data is generated using a spatial model. Similarly, Lepelley et al. [21] and Miller [29] estimate the frequency with which IRV produces monotonicity paradoxes under various theoretical models of voter behavior. Lepelley and Merlin [22] analyze the frequency with which IRV demonstrates a no-show paradox in three-candidate elections under the impartial culture and impartial anonymous culture models. Kamwa et al. [18] analyze the probability that IRV (among other methods) demonstrates no-show or negative participation paradoxes on single-peaked domains.
The third strand is the expanding empirical social choice literature on IRV, enabled by the increasing availability of ranked-choice ballot data from jurisdictions such as Australia [43], the United States [34], and Scotland [24]. Some of this literature concerns the empirical probability that IRV produces a particular kind of paradox. Graham-Squire and Zayatz [15], Graham-Squire and McCune [17], and McCune and Graham-Squire [24] estimate the frequency with which real-world IRV elections demonstrate monotonicity and no-show paradoxes. Other empirical work includes case studies of particularly interesting elections which demonstrate a range of different paradoxes; see [20, 25], for example. Beyond work on paradoxes, empirical research on IRV has analyzed rates of ballot errors in IRV elections [7, 35], the extent of ballot exhaustion [6, 17], the rate at which IRV exhibits the so-called “spoiler effect” [27], the extent to which ballot truncation can affect the IRV winner [9, 19, 44], and the frequency with which IRV selects the Condorcet winner [17, 42, 43]. There is also a growing, related empirical literature concerning the method of the single transferable vote (STV), a multiwinner version of IRV which is designed to achieve proportional representation given a set of preference ballots. This literature also investigates paradoxical and Condorcet-related outcomes [24, 26], as well as STV’s propensity to achieve proportional representation [1, 3, 30].
Previous work on the reinforcement paradox is fairly limited. The notion of a reinforcement paradox dates to the work of Young [45, 46], who defined the concept and proved essentially that the only reasonable voting methods not susceptible to the paradox are positional scoring rules. Felsenthal and Nurmi [13] show that various Condorcet inconsistent voting rules (including IRV) are susceptible to the paradox even in “tame” profiles in which a Condorcet winner exists and the voting rule chooses this candidate. Brandt et al. [2] provide some “minimal” examples of the reinforcement paradox under IRV where the number of candidates and voters is small. Brandt et al. [4] analyze the frequency of reinforcement paradoxes under various Condorcet consistent voting rules. We are aware of only two previous papers which attempt to measure the frequency with which IRV produces reinforcement paradoxes, both of which focus on the three-candidate case. Plassmann and Tideman [37] measure this frequency under a spatial model, and Courtin et al. [8] use the impartial culture and impartial anonymous culture models of voter behavior. Our work differs from these studies in significant aspects which we explicate more fully below. In brief, these two studies tend to generate two elections and then say that the reinforcement paradox has occurred if these two elections have the same winner but when the ballots from the two elections are combined the winner is different. By contrast, we consider only one election and say the paradox has occurred in this election if there exists a way to partition the ballots into two sub-elections where the winners of the sub-elections are the same but differ from the winner of the original election.
Of course, there is a much broader literature concerning paradoxes and voting methods other than IRV, and we do not attempt to survey that literature here. See [11, 31, 36] for an examination of the susceptibility of Condorcet consistent methods to various types of no-show paradox, for example.
The paper is structured as follows. Section 2 gives preliminaries, including formal definitions of IRV and the reinforcement paradox. In Section 3 we provide necessary and sufficient conditions for a three-candidate election to demonstrate a reinforcement paradox and discuss some of the consequences of these conditions. Proposition 1 gives these conditions assuming all voters provide a complete ranking of the candidates while Proposition 2 gives conditions when some voters cast partial ballots. Section 4 considers the minimum percentage a losing candidate can earn and yet the election demonstrates a reinforcement paradox. Section 5 estimates the probability that an election demonstrates a paradox under various theoretical models of voter behavior. Section 6 analyzes real-world IRV elections from the United State and Scotland. This section provides empirical probabilities using a large dataset of elections, and provides a probabilistic analysis using non-parametric bootstrapping. Section 7 gives a discussion of our results, and Section 8 concludes.
2. Preliminaries
In this section we define IRV and types of reinforcement paradoxes, accompanied by examples.
In an election which uses IRV voters cast preference ballots with a (possibly partial) linear ranking of the candidates. After an election, the ballots are aggregated into a preference profile, which shows the number of each type of ballot cast. For example, Table 1 shows a preference profile involving the three candidates $C$, $G$, and $H$. The number 2211 denotes that 2211 voters ranked $G$ first, $H$ second, and $C$ third. In the real-world IRV elections we analyze in Section 6, voters are allowed to provide partial rankings. Table 1 has three columns where voters do not provide a complete ranking–2003 voters rank only $G$ on their ballots, for example.
An IRV election proceeds in rounds where, in each round, the number of first-place votes for each candidate is calculated. If a candidate has a majority of the (remaining) first-place votes, they are declared the winner. If no candidate receives a majority of votes, the candidate with the fewest first-place votes is eliminated and the names of the remaining candidates on the affected ballots are moved up. The process repeats until a candidate is declared the winner. [^1] The issue of ties does not arise in our work in this paper, and thus we ignore the possibility of multiple winners. In the case when partial ballots are allowed, any ballot in which all candidates have been eliminated is considered “exhausted” and plays no further role in the determination of the winner. [^2] In theory there are many ways we could treat exhausted ballots besides removing them from the election. However, as far as we are aware, any actual election office treats exhausted ballots in this fashion.
When there are only three candidates, at most two rounds are required to determine a winner; in general, several rounds may be required. We focus only on the three-candidate case and thus do not consider more than two rounds of ballot counting. The following example demonstrates the IRV algorithm using the 2022 Alaska Senate District E election in Alaska.
Num. Voters 2211 1439 2003 2605 241 2689 2230 418 2873
1st choice $G$ $G$ $G$ $H$ $H$ $H$ $C$ $C$ $C$
2nd choice $H$ $C$ $G$ $C$ $G$ $H$
3rd choice $C$ $H$ $C$ $G$ $H$ $G$
Table 1. Vote totals in the 2022 Senate District E election in Alaska.
Example 1.
(2022 Alaska Senate District E.) This election contained the three candidates Roselynn Cacy (C), Cathy Giessel (G), and Roger Holland (H). The preference profile is displayed in Table 1.333Our vote totals differ from the officially reported vote totals by up to three votes. This minor discrepancy could be due to a recount which occurred after the election, or because the Alaska Division of Elections has not been completely clear about how it processes “edge case” ballots. Our analysis is unaffected.
Initially, the vote totals for Cacy, Giessel, and Holland are 5521, 5653, and 5535, respectively. Cacy has the fewest votes and is eliminated from the election. As a result, 2230 votes are transferred to Giessel and 418 to Holland, and Giessel wins in the last round 7883 votes to 5953.
While Giessel is the winner, this election has the strange property that Cacy can point to two disjoint groups of voters such that Cacy would win under IRV if we use only the ballots from each group. Similarly, Holland can also break the electorate into two disjoint groups such that Holland wins when using the ballots from each group separately. We illustrate this paradoxical behavior in the next example.
Example 2.
In the 2022 Senate District E election in Alaska, suppose we partition the ballots into two ballot sets $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ as shown for Partition 1 in Table 2. If we use just the ballots from $\mathcal{B}_{1}$, $H$ wins under IRV, as $C$ is eliminated first and enough votes are transferred to $H$ for him to win against $G$ win in the final round. Similarly, if we use only the ballots from $\mathcal{B}_{2}$ then $H$ also wins under IRV, as in this case $G$ is eliminated first and enough votes are transferred to $H$ for him to win against $C$ in the final round. Thus, Roger Holland is able to point to two disjoint sub-electorates in which he would win the election using just the ballots from each sub-electorate, yet when we merge the ballots from the sub-electorates to recover the original ballot data, Holland loses.
Table 2 provides a second partition of the ballots, Partition 2, in which $C$ wins under IRV when using only the ballots from $\mathcal{B}_{1}$ and only the ballots from $\mathcal{B}_{2}$, yet when we merge the ballots to recover the original election data, $C$ loses.
For both Cacy and Holland, in this election win $+$ win $=$ loss in some sense.
Num. Voters 2211 1439 2003 2605 241 2689 2230 418 2873
1st choice $G$ $G$ $G$ $H$ $H$ $H$ $C$ $C$ $C$
2nd choice $H$ $C$ $G$ $C$ $G$ $H$
3rd choice $C$ $H$ $C$ $G$ $H$ $G$
Partition 1 $\mathcal{B}_{1}$ 0 1439 1500 0 0 2689 0 418 2200
$\mathcal{B}_{2}$ 2211 0 503 2605 241 0 2230 0 673
Partition 2 $\mathcal{B}_{1}$ 0 1439 0 2605 0 0 0 0 1440
$\mathcal{B}_{2}$ 2211 0 2003 0 241 2689 2230 418 1433
Table 2. Two different partitions of the ballots in the 2022 Senate District E election in Alaska.
Example 2 demonstrates the notion of a reinforcement paradox, the subject of this article. The study of this paradox dates to the work of Young [45]. Young considered a voting method as a function $f$ which takes as input a set of ranked ballots and outputs an election winner (or a set of winners, if there is some sort of tie). If $\mathcal{B}$ is a set of ballots then $f(\mathcal{B})$ denotes the election winner under a voting method $f$, which leads to the definition of consistency found in [45].
Definition 1.
Let $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ be two sets of ballots over the same candidate set. A voting method $f$ is said to be consistent if whenever $f(\mathcal{B}_{1})=f(\mathcal{B}_{2})$ then $f(\mathcal{B}_{1}\cup\mathcal{B}_{2})=f(\mathcal{B}_{1})$.
A voting method that is not consistent is said to be susceptible to the reinforcement paradox. This paradox has other names, such as the “multiple districts paradox” [14], and is a social choice analogue of the classical Simpson’s Paradox from statistics.
Definition 2.
An election with ballot set $\mathcal{B}$ demonstrates a reinforcement paradox under a voting method $f$ if there exists a partition of the ballot set $\mathcal{B}=\mathcal{B}_{1}\cup\mathcal{B}_{2}$ such that $f(\mathcal{B}_{1})=f(\mathcal{B}_{2})$ and $f(\mathcal{B}_{1})\neq f(\mathcal{B}_{1}\cup\mathcal{B}_{2})$.
In [46], Young showed that consistency and a few other uncontroversial axioms completely characterize positional scoring rules. Thus, any “reasonable” voting method which is not a scoring rule can demonstrate a reinforcement paradox for some election. We analyze only IRV, and thus the function $f$ represents that voting method throughout the article.
We conclude this section by providing additional notation for preference profiles and giving vocabulary for different types of reinforcement paradox.
Assume the preference profiles for complete and partial rankings are as shown in Tables 3 and 4 where in each case the numbers $X_{i},Y_{i}$, and $Z_{i}$ indicate the number of voters with the corresponding ranking and that these numbers sum to the total number of voters, which we denote $V.$ Note in Table 4 we do not list ballots of length two, as such ballots convey the same ranking information as the corresponding complete ballots.
Num. Voters $X_{1}$ $X_{2}$ $Y_{1}$ $Y_{2}$ $Z_{1}$ $Z_{2}$
1st choice $A$ $A$ $B$ $B$ $C$ $C$
2nd choice $B$ $C$ $A$ $C$ $A$ $B$
3rd choice $C$ $B$ $C$ $A$ $B$ $A$
Table 3. Preference profile with complete preferences.
Num. Voters $X_{1}$ $X_{2}$ $X_{3}$ $Y_{1}$ $Y_{2}$ $Y_{3}$ $Z_{1}$ $Z_{2}$ $Z_{3}$
1st choice $A$ $A$ $A$ $B$ $B$ $B$ $C$ $C$ $C$
2nd choice $B$ $C$ $A$ $C$ $A$ $B$
3rd choice $C$ $B$ $C$ $A$ $B$ $A$
Table 4. Preference profile with some partial preferences.
When there are three candidates, we refer to the candidate with the smallest number of first-place votes as the plurality loser. This candidate can never be the winner in an IRV election since they are eliminated in the first round. We refer to the remaining losing candidate, (who may receive either the largest or second-largest number of first-place votes) as the runner-up.
Definition 3.
An election demonstrates a plurality loser reinforcement paradox if there exists a partition of the ballots $\mathcal{B}=\mathcal{B}_{1}\cup\mathcal{B}_{2}$ such that the plurality loser is the IRV winner when using only the ballots in $\mathcal{B}_{1}$ or only the ballots in $\mathcal{B}_{2}$.
Definition 4.
An election demonstrates a runner-up reinforcement paradox if there exists a partition of the ballots $\mathcal{B}=\mathcal{B}_{1}\cup\mathcal{B}_{2}$ such that the runner-up is the IRV winner when using only the ballots in $\mathcal{B}_{1}$ or only the ballots in $\mathcal{B}_{2}$.
If an election demonstrates both a plurality loser and runner-up paradox, we say the election demonstrates a double reinforcement paradox.
We now turn to the problem of determining if a ballot set $\mathcal{B}$ in a three-candidate election can be partitioned to produce a runner-up or plurality loser reinforcement paradox.
3. Necessary and Sufficient Conditions
The purpose of this article is to investigate the probability with which a three-candidate election demonstrates a reinforcement paradox. To examine such probabilities, we first must know when an election demonstrates a paradox. This section characterizes when an election demonstrates a reinforcement paradox, identifying sets of inequalities in terms of the parameters provided in Table 3 for the case of complete ballots or the parameters provided in Table 4 for the partial ballot case.
First, we observe that if one of the candidates earns an initial majority of first-place votes (we term such a candidate a “majority candidate”) then the election cannot demonstrate a reinforcement paradox. That is, we cannot observe a paradox in the complete ballot case if $X_{1}+X_{2}>Y_{1}+Y_{2}+Z_{1}+Z_{2}$ or $Y_{1}+Y_{2}>X_{1}+X_{2}+Z_{1}+Z_{2}$ or $Z_{1}+Z_{2}>X_{1}+X_{2}+Y_{1}+Y_{2}$; similarly for the partial ballot case. The reason is that if a majority candidate exists, that candidate will earn a majority of first-place votes in either $\mathcal{B}_{1}$ or $\mathcal{B}_{2}$ for any partition of the ballot set.
Observation 1.
If an election contains a majority candidate then the election cannot demonstrate a reinforcement paradox.
In the complete ballot case, let $X=X_{1}+X_{2}$, $Y=Y_{1}+Y_{2}$, and $Z=Z_{1}+Z_{2}$; in the partial ballot case, let $X^{\prime}=X_{1}+X_{2}+X_{3}$, $Y^{\prime}=Y_{1}+Y_{2}+Y_{3}$, and $Z^{\prime}=Z_{1}+Z_{2}+Z_{3}$. We assume without loss of generality that $A$ is the IRV winner and $C$ is the plurality loser. Thus, in the complete ballot case we assume
$Z<X,\ Z<Y\text{ and }Y+Z_{2}<X+Z_{1}.$
In the partial ballot case we assume
$Z^{\prime}<X^{\prime},\ Z^{\prime}<Y^{\prime}\text{ and }Y^{\prime}+Z_{2}<X^{\prime}+Z_{1}.$
Under these assumptions, the following propositions characterize when a three-candidate election demonstrates a reinforcement paradox of either type.
Proposition 1.
(Complete Ballots). An election demonstrates a runner-up reinforcement paradox if and only if
  1. (1)
    $Y>X\text{ and }Y>-X_{1}+X_{2}+Z$ or
  2. (2)
    $Y>X-Z_{2}\text{ and }Y>-X_{1}+X_{2}+Z_{1}-Z_{2}.$
If the election satisfies (1) then the runner-up can win a majority in one of the sub-elections and if the election satisfies (2) then the runner-up does not win a majority in either sub-election.
An election demonstrates a plurality loser reinforcement paradox if and only if
$Z>X-Y_{2},\ Z>X_{1}-X_{2}+Y_{1}-Y_{2},\ Z>-X_{2}+Y,\text{ and }\ Z>\frac{1}{3}(X+Y).$
Proposition 2.
(Partial Ballots). An election demonstrates a runner-up reinforcement paradox if and only if
  1. (1)
    $Y^{\prime}>X^{\prime}\text{ and }Y^{\prime}>-X_{1}+X_{2}+Z^{\prime},$ or
  2. (2)
    $Y^{\prime}>X^{\prime}-Z_{2},Y^{\prime}>-X_{1}+X_{2}+Z_{1}-Z_{2}$, and $Y^{\prime}>\frac{1}{3}(-X_{1}+2X_{2}+2X_{3}+Z_{1}-Z_{2}).$
If the election satisfies (1) then runner-up can win a majority in one of the sub-elections and if the election satisfies (2) then the runner-up does not win a majority in either sub-election.
An election demonstrates a plurality loser reinforcement paradox if and only if
$Z^{\prime}>X^{\prime}-Y_{2},\ Z^{\prime}>X_{1}-X_{2}+Y_{1}-Y_{2},\ Z^{\prime}>-X_{2}+Y^{\prime},\ Z^{\prime}>\frac{1}{3}(X^{\prime}+Y^{\prime}).$
The conditions in Proposition 2 reduce to those in Proposition 1 when $X_{3}=Y_{3}=Z_{3}=0$ for both runner-up and plurality loser reinforcement paradoxes. The third condition $Y^{\prime}>\frac{1}{3}(-X_{1}+2X_{2}+2X_{3}+Z_{1}-Z_{2})$ in Case (2) of Proposition 2 reduces to $Y>\frac{1}{3}(-X_{1}+2X_{2}+Z_{1}-Z_{2})$ and does not have an analogue in Proposition 1 because it is implied by the other conditions in Case (2). To see this, note that
$X_{1}-2X_{2}$ $+3Y-Z_{1}+Z_{2}=(X-Y+Z_{1}-Z_{2})$
$+2(-X+Y+Z_{2})+(X_{1}-X_{2}+Y-Z_{1}+Z_{2})+X_{1}+(Y-Z).$
But $-X+Y+Z_{2}>0$ and $X_{1}-X_{2}+Y-Z_{1}+Z_{2}>0$ by the remaining two conditions in Case (2) and $Y-X>0$ by assumption. Hence $X_{1}-2X_{2}+3Y-Z_{1}+Z_{2}>0$, which implies $Y>\frac{1}{3}(-X_{1}+2X_{2}+Z_{1}-Z_{2})$.
Before sketching the proofs of the propositions, we make a few observations. Firstly, a runner-up and plurality loser reinforcement paradoxes can occur when the IRV and plurality winners are distinct or when they coincide, assuming either complete or partial ballots. Secondly, in Case (2) of the runner-up paradox and in the plurality loser paradox it is easily seen that the runner-up (plurality loser respectively) wins both sub-elections only when different candidates are eliminated in the two sub-elections.
Thirdly, in Case (1) of the runner-up paradox, the runner-up must be the plurality winner, and the proof of the propositions show that if the two conditions are met, the runner-up wins a majority of first-place votes in one sub-election and wins in two rounds with the IRV winner being the plurality loser in the other sub-election. Note in Case (1), if $X_{1}>X_{2}$ then $Y^{\prime}>-X_{1}+X_{2}+Z^{\prime}$ is automatically satisfied. Thus, if among the voters whose first choice is the IRV winner, more voters second-rank the runner-up than second-rank the plurality loser, the paradox occurs whenever the runner up is the plurality winner ($Y^{\prime}>X^{\prime}$ or $Y>X$ in the complete ballot case).
We sketch the proof of Proposition 1 for the plurality loser. The remaining proofs are similar.
Proof.
Let $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ be a partition of the original election into two sets of ballots in which $C$ is the IRV winner. Let $\mathcal{B}_{1}$ have voter profile as shown in Table 3 where $0\leq x_{1}\leq X_{1}$ replaces $X_{1}$, $0\leq x_{2}\leq X_{2}$ replaces $X_{2}$ and so on.
Candidate $C$ cannot have a majority of first-place votes in both sub-elections (otherwise it would have a majority in the original election). We claim that $C$ cannot have a majority of first-place votes in either sub-election. To see this, suppose $C$ has a majority in $\mathcal{B}_{1}$. If $A$ is the plurality loser in $\mathcal{B}_{2}$, then
$x_{1}+x_{2}<z_{1}+z_{2}-y_{1}-y_{2}\mbox{ and }X_{1}+X_{2}-x_{1}-x_{2}<Z_{1}+Z_{1}-z_{1}-z_{2}.$
Adding these inequalities yields $X_{1}+X_{2}<Z_{1}+Z_{2}-y_{1}-y_{2}<Z_{1}-Z_{2}$, which is not the case. The same contradiction arises if $B$ is the plurality loser in $\mathcal{B}_{2}.$ This proves the claim.
Thus, $C$ must win in the second round after the elimination of a plurality loser in both sub-elections. The two sub-elections cannot have the same plurality loser (otherwise that candidate would be a plurality loser when the ballots were combined). So assume $B$ is the plurality loser in $\mathcal{B}_{1}$ and $A$ is the plurality loser in $\mathcal{B}_{2}.$ This yields the following set of inequalities.
(1) $y_{1}+y_{2}$ $<x_{1}+x_{2}$
(2) $y_{1}+y_{2}$ $<z_{1}+z_{2}$
(3) $x_{1}+x_{2}+y_{1}$ $<z_{1}+z_{2}+y_{2}$
(4) $X_{1}+X_{2}-x_{1}-x_{2}$ $<Y_{1}+Y_{2}-y_{1}-y_{2}$
(5) $X_{1}+X_{2}-x_{1}-x_{2}$ $<Z_{1}+Z_{2}-z_{1}-z_{2}$
(6) $Y_{1}+Y_{2}-y_{1}-y_{2}+X_{1}-x_{1}$ $<Z_{1}+Z_{2}-z_{1}-z_{2}+X_{2}-x_{2}$
Note that since $Z_{1}+Z_{2}<Y_{1}+Y_{2}$, Equations (1) and (5) imply
$X_{1}+X_{2}-x_{1}-x_{2}<Z_{1}+Z_{2}-z_{1}-z_{2}<Y_{1}-Y_{2}-y_{1}-y_{2}.$
So we can eliminate (4). Recalling that $z=z_{1}+z_{2}$ and $Z=z_{1}+z_{2},$ and incorporating the conditions $0\leq y_{1}\leq Y_{1}$, etc. into the constraints,we can rewrite these inequalities as
$\max\{0,$ $X_{1}-X_{2}+Y_{1}+Y_{2}-Z-x_{1}+x_{2}-y_{2}+z\}<y_{1}<\min\{Y_{1},$
$x_{1}+x_{2}-y_{2},z-y_{2},-x_{1}-x_{2}+z+y_{2}\}$
and
$0<z<\min\{Z,-X_{1}-X_{2}+Z+x_{1}+x_{2}\}$
Since every lower bound of $y_{1}$ must be bounded above by every upper bound of $y_{1},$ these bounds on $y_{1}$ lead to the following conditions
00 0 $<x_{1}+x_{2}-y_{2}$
$\quad\Rightarrow y_{2}<x_{1}+x_{2}$
00 0 $<z-y_{2}$
$\quad\Rightarrow y_{2}<z$
00 0 $<-x_{1}-x_{2}+z+y_{2}$
$\quad\Rightarrow x_{1}+x_{2}-y_{2}<z$
$X_{1}-X_{2}+Y_{1}+Y_{2}-Z-x_{1}+x_{2}-y_{2}+z$ $<Y_{1}$
$\quad\Rightarrow z<-X_{1}+X_{2}-Y_{2}+Z+x_{1}-x_{2}+y_{2}$
$X_{1}-X_{2}+Y_{1}+Y_{2}-Z-x_{1}+x_{2}-y_{2}+z$ $<x_{1}+x_{2}-y_{2}$
$\quad\Rightarrow z<-X_{1}+X_{2}-Y_{1}-Y_{2}+Z+2x_{1}$
$X_{1}-X_{2}+Y_{1}+Y_{2}-Z-x_{1}+x_{2}-y_{2}+z$ $<z-y_{2}$
$\quad\Rightarrow X_{1}-X_{2}+Y_{1}+Y_{2}-Z+x_{2}<x_{1}$
$X_{1}-X_{2}+Y_{1}+Y_{2}-Z-x_{1}+x_{2}-y_{2}+z$ $<-x_{1}-x_{2}+z+y_{2}$
$\quad\Rightarrow\frac{1}{2}(X_{1}-X_{2}+Y_{1}+Y_{2}-Z)+x_{2}<y_{2}$
Adding the new conditions on $z$ into the bounds on $z$, we have
$\max\{0,\,$ $y_{2},\,x_{1}+x_{2}-y_{2}\}<z<\min\{Z,-X_{1}-X_{2}+Z+x_{1}+x_{2},$
$-X_{1}+X_{2}-Y_{2}+Z+x_{1}-x_{2}+y_{2},-X_{1}+X_{2}-Y_{1}-Y_{2}+Z+2x_{1},\}$
These conditions, along with the other conditions above in which $z$ is eliminated, only involve $z,y_{2},x_{1},x_{2}$ and the parameters of the original equation $X_{1},X_{2}$, etc. Requiring each lower bound of $z$ to be bounded above by every upper bound of $z$ as we did for $y_{1}$, creates another set of inequalities involving $y_{2},x_{1},x_{2}$ and the parameters $X_{1},X_{2}$, etc. We continue in this way, eliminating the variables $y_{2},x_{1}$ and $x_{2}$ in turn until the remaining conditions involve only $X_{1},X_{2}$, etc. Simplified, these reduce to the inequalities in Proposition 1. ∎
The inequalities created in each step of the iterative process in the proofs of the propositions can be used in reverse to identify election partitions leading to the paradoxes. For instance, from the proof of Case (1) of Proposition 1, we can derive the following bounds for the sub-elections.
  • $\max\{X_{2}+X_{3}-Y^{\prime}+Z^{\prime},\frac{1}{2}(X^{\prime}-Y^{\prime}+Z^{\prime})\}<y^{\prime}<\min\{-X^{\prime}+Y^{\prime}+Z^{\prime}\}$
  • $\max\{X^{\prime}-Y^{\prime}+Z^{\prime},\frac{1}{2}(X^{\prime}-Y^{\prime}+Z^{\prime})+\frac{1}{2}y^{\prime},-Y^{\prime}+Z^{\prime}+y^{\prime}\}<z^{\prime}<\min\{Z^{\prime},X_{1}+y^{\prime},2y^{\prime}\}$
  • $\max\{0,X_{2}-Y^{\prime}+Z^{\prime}+y^{\prime}-z^{\prime}\}<x_{2}<\min\{X_{2},z^{\prime},X_{1}+y^{\prime}-z^{\prime},y^{\prime}-\frac{1}{2}z^{\prime},\frac{1}{2}y^{\prime}\}$
  • $\max\{0,-y^{\prime}+z^{\prime}+x_{2},X_{1}+X_{2}-Y^{\prime}+Z^{\prime}-x_{2}+y^{\prime}-z^{\prime}\}<x_{1}<\min\{X_{1},y^{\prime}-x_{2},z^{\prime}-x_{2}\}$
  • $\max\{0,X^{\prime}-Y^{\prime}+Z^{\prime}-x_{1}-x_{2}+y^{\prime}-z^{\prime}\}<x_{3}<\min\{X_{3},y^{\prime}-x_{1}-x_{2},z^{\prime}-x_{1}-x_{2}\}$
Care must be taken when applying these bounds: the inequalities hold only in the limit as no allowance is made for the discrete mature of ballot counts and the possibility of ties. However, if the $X_{1},X_{2},$ etc. values in the original election are sufficiently distinct, these bounds can be used to identify two sub-elections.
For instance, in the complete ballot case, if $X_{1}=100,X_{2}=105,Y=215,Z_{1}=110$ and $Z_{2}=90$, these conditions allow us to select (among other possibilities)
$x_{1}=99,x_{2}=1,y=106\mbox{ and }z=197,$
which leads to
$X_{1}-x_{1}=1,X_{2}-x_{2}=104,Y-y=109,Z-z=3.$
It is easily seen that $A$ wins the original election, but in $\mathcal{B}_{1}$, candidate $B$ wins in two rounds after $A$ is eliminated and in $\mathcal{B}_{2}$, candidate $B$ has a majority.
3.1. Consequences of the Algebraic Conditions
The inequalities identified in Propositions 1 and 2 allow us to identify further anomalies. For instance, Table 5 demonstrates that it is possible for a candidate to be the IRV winner, a Condorcet winner and a plurality winner and yet the profile still satisfies the requirements in Propositions 1 for both a runner-up and plurality loser paradox.
Num. Voters 1001 50 521 480 561 439
1st $A$ $A$ $B$ $B$ $C$ $C$
2nd $B$ $C$ $A$ $C$ $A$ $B$
3rd $C$ $B$ $C$ $A$ $B$ $A$
Table 5. (Top) A profile with the IRV, plurality, and Condorcet winners coinciding that demonstrates both runner-up and plurality loser paradoxes.
The next example shows that reinforcement paradoxes can be nested: an election may exhibit a runner-up or plurality loser paradox and each sub-election demonstrate an additional reinforcement paradoxes.
Example 3.
Consider the profile in Table 6 where $A$ is the IRV winner. The election demonstrates a runner-up paradox and $B$ wins in both sub-elections $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$. However each of these sub-elections themselves satisfies the conditions for a runner-up paradox. This means that sub-election $\mathcal{B}_{1}$, in which $C$ is the plurality loser and $A$ is the runner-up, can be further subdivided into two smaller sub-elections in which $A$ is the IRV winner. (One such partition is $\mathcal{B}_{11}$ which corresponds to $X_{1}=12$, $X_{2}=31$, $Y_{1}=0$, $Y_{2}=25$, $Z_{1}=0$ and $Z_{2}=17$ and $\mathcal{B}_{12}$ which corresponds to $X_{1}=28$, $X_{2}=31$, $Y_{1}=5$,$Y_{2}=26$, $Z_{1}=1$ and $Z_{2}=31$.) And sub-election $\mathcal{B}_{2}$, in which $A$ is the plurality loser and $C$ is the runner-up, can be further subdivided into two smaller sub-elections in which $C$ is the IRV winner. (One such partition is $\mathcal{B}_{21}$ which corresponds to $X_{1}=50$, $X_{2}=0$, $Y_{1}=4$, $Y_{2}=25$, $Z_{1}=30$ and $Z_{2}=0$ and $\mathcal{B}_{22}$ which corresponds to $X_{1}=10$, $X_{2}=38$, $Y_{1}=41$, $Y_{2}=54$, $Z_{1}=69$ and $Z_{2}=2$.)
Num. Voters 100 100 50 130 100 50
1st choice $A$ $A$ $B$ $B$ $C$ $C$
2nd choice $B$ $C$ $A$ $C$ $A$ $B$
3rd choice $C$ $B$ $C$ $A$ $B$ $A$
Partition $\mathcal{B}_{1}$ 40 62 5 51 1 48
$\mathcal{B}_{2}$ 60 38 45 79 99 2
Table 6. A profile that demonstrates a runner-up paradox with sub-elections that demonstrate additional reinforcement paradoxes.
Additional instances of the paradox may occur in which an election can be divided into three or more sub-elections in which either the runner-up or plurality loser wins.
Num. Voters 1000 999 560 438
1st choice $A$ $B$ $C$ $C$
2nd choice $C$ $C$ $A$ $B$
3rd choice $B$ $A$ $B$ $A$
Table 7. A profile that demonstrates a plurality loser, but not a runner-up, reinforcement paradox.
Finally, we note that elections may be susceptible to a plurality loser paradox but not a runn-erup paradox, as demonstrated by the profile in Table 7 where neither Condition (1) nor (2) of Proposition 1 is satisfied, but the plurality loser inequalities are satisfied. Such preference profiles can arise when voter preferences are “single-peaked.” In this case, candidate $C$ is like a “centrist” candidate in a spatial model, as if $A$, $B$, and $C$ have been arranged on a left-to-right line, with $C$ in the middle.
Of course, real-world elections typically do not return ballot data resembling the profile in Table 7. In fact as we will see in Section 6 in real elections, the IRV winner tends to be a much stronger candidate than the plurality loser. Thus we expect that if an election demonstrates a plurality loser reinforcement paradox in this data then the election will also demonstrate a runner-up paradox.
4. Threshold Results
In this section we examine the minimal percentage each candidate can earn while the election still demonstrates a reinforcement paradox. For some classical paradoxes such as upward and downward monotonicity paradoxes, an election can exhibit such a paradox only if each of the three candidates is not “weak.” We show this is not the case for a general reinforcement paradox in Proposition 3, where the plurality loser can have an arbitrarily low percentage of the first-places votes and yet an election can demonstrate a reinforcement paradox. We show in Proposition 4 that the candidate affected by the paradox must earn at least 25% of the first place vote, so that if an election demonstrates a runner-up (respectively plurality loser) reinforcement paradox then the IRV runner-up (respectively plurality loser) must earn at least 25% of the first-place votes.
In this section we assume that the number of voters approaches infinity, so that we use proportions of votes rather than integers. We also assume that each voter provides a complete ranking as this does not affect the logic of the proofs.
Proposition 3.
The plurality loser can have arbitrarily small first-place support and yet the election demonstrate the reinforcement paradox.
Proof.
Consider the profile in Table 8, where $p_{A}$ is the proportion of voters who cast the ballot $A\succ B\succ C$ and also the proportion of voters who cast the ballot $A\succ C\succ B$, and $\epsilon$ and $\delta$ are positive and very small. Furthermore, assume $0.52p_{A}+\epsilon<$ 0.5 and $14p_{A}+\epsilon+\delta=$ 1 (which implies $\delta>\epsilon)$. Note that $A$ is the IRV winner and $C$ is the plurality loser.
This profile satisfies condition (2) of Proposition 1 and thus the election demonstrates a runner-up reinforcement paradox. For completeness, we include a partition of the electorate into two sub-electorates, as shown in the table. Let $\delta_{B}$ and $\delta_{A}$ be positive small parameters such that $20<\delta_{B}<\delta_{A}<\delta/$ 2. Then $B$ wins when using the ballots of both sub-electorates separately. Since $\delta$ is arbitrary, the proposition is proven.∎
Prop. Voters $p_{A}$ $p_{A}$ $2p_{A}+\epsilon$ $\delta$
1st choice $A$ $A$ $B$ $C$
2nd choice $B$ $C$ $A$ $A$
3rd choice $C$ $B$ $C$ $B$
Partition $\mathcal{B}_{1}$ $\delta-\delta_{A}$ 0 $\delta-\delta_{B}$ $\delta$
$\mathcal{B}_{2}$ $p_{A}-\delta+\delta_{A}$ $p_{A}$ $2p_{A}+\epsilon-\delta+\delta_{B}$ 0
Table 8. A profile in which the plurality loser earns a small proportion of first-place votes and the election demonstrates a (runner-up) reinforcement paradox.
Thus, for a large enough electorate the plurality loser can control an arbitrarily small amount of first-place votes and the election can demonstrate a reinforcement paradox. This proposition distinguishes reinforcement paradoxes from other classical paradoxes such as upward or downward monotonicity paradoxes. For a three-candidate election to demonstrate an upward (resp. downward) paradox, each candidate must earn at least 25% (respectively $16\frac{2}{3}\$) of the first-place votes [29]. That is, for some classical paradoxes none of the candidates can have an extremely weak first-place showing; each candidate must be “viable” in some sense. Proposition 3 shows this is not the case for the reinforcement paradox. However, in the next proposition we show that for a runner-up (resp. plurality) reinforcement paradox, the runner-up (resp. plurality loser) must earn at least 25% of the first-place votes.
Proposition 4.
The IRV runner-up (resp. plurality loser) can have, in the limit, as little as 25% of the first-place vote and the election demonstrate a runner-up (resp. plurality loser) reinforcement paradox. In addition, the IRV winner can have, in the limit, as much as 50% of the first-place vote and the election demonstrate either a runner-up or plurality loser reinforcement paradox.
Proof.
Runner-up case. Recall that in order for an election to demonstrate a reinforcement paradox, no candidate can earn an initial majority, and thus candidate $A$ must earn less than 50% of the first-place votes. This implies $B$ must earn at least $25\%$ of the first-place vote.
To see that that bounds are tight, consider the profile in Table 9 where $\epsilon_{A},\epsilon_{B},\epsilon_{1}$ and $\epsilon_{2}>0$ are arbitrarily small parameters. If $\epsilon_{B}>\epsilon_{2}-\epsilon_{A}$ then this profile satisfies the conditions in (2) of Proposition 1.
Num. Voters $0.50-\epsilon_{A}$ $0.25+\epsilon_{B}$ $\epsilon_{1}$ $0.25-\epsilon_{2}$
1st choice $A$ $B$ $C$ $C$
2nd choice $B$ $A$ $A$ $B$
3rd choice $C$ $C$ $B$ $A$
Table 9. Minimum profile for runner-up reinforcement paradox
Plurality loser case. From Proposition 1, $Z>\frac{1}{3}(X+Y)$ and thus the plurality loser must have at least $25\%$ of the first-place vote. On the other hand, consider the profile in Table 10 in which the plurality loser has just over 25% of the vote and yet, for some suitably small $\epsilon>0,$ satisfies the plurality loser inequalities from Proposition 1. ∎
Num. Voters $0.375-\epsilon$ $0.375-2\epsilon$ $0.25+3\epsilon$
1st choice $A$ $B$ $C$
2nd choice $C$ $C$ $A$
3rd choice $B$ $A$ $B$
Table 10. Minimum profile for plurality loser reinforcement paradox
We conclude this section by noting that there are elections that demonstrate a reinforcement paradox in which one of the sub-elections is arbitrarily small.
Example 4.
Let $m\in\mathbb{N}$ be a large even number. Consider an election with ballot set $\mathcal{B}$ where $m$ ballots have the form $A\succ B$, $m+1$ have the form $B\succ A$, $m/2$ have the form $C\succ A$, and $m/2-2$ have the form $C\succ B$. Then $A$ wins this election, since initially they have one fewer vote than $B$ but the elimination of $C$ allows $A$ to win by one vote.
Let $\mathcal{B}_{1}$ consist of two ballots of the form $A\succ B$, three of the form $B\succ A$, and four of the form $C\succ A$. Let $\mathcal{B}_{2}$ consist of $m-2$ ballots of the form $A\succ B$, $m-2$ of the form $B\succ A$, $m/2-4$ of the form $C\succ A$, and $m/2-2$ of the form $C\succ B$. Then $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ form a partition of $\mathcal{B}$, and candidate $B$ wins both of these sub-elections. Since $m$ can be arbitrarily large, this implies there exists an election demonstrating a reinforcement paradox where one of the sub-elections can be arbitrarily small.
Example 4 is interesting as it suggests that $B$ was “close” to being a winner in the original election. In fact, $B$ might feel it particularly unfair that the addition of a handful of ballots (from $\mathcal{B}_{1}$, where $B$ was a winner) caused $B$ to go from winner (in $\mathcal{B}_{2}$) to loser (in $\mathcal{B}_{1}\cup\mathcal{B}_{2}$).
5. Results under various models
In this section we use different models of voter behavior to generate millions of synthetic elections. Among the many possible models we could use, we choose models which have been used extensively in prior social choice research. For each model, we generate 100,000 profiles and check whether a generated election demonstrates a reinforcement paradox. Thus, our primary tool for providing (estimated) probabilities is Monte Carlo simulation, although we can provide exact probabilities using other means for the impartial anonymous culture (IAC) model.
We are interested in probabilities for large electorates, as the real-world elections we analyze in the next section generally have thousands of voters. Thus, our simulations use an electorate size of $V=1001$ voters. We also ran simulations with $V\in\{601,2003,3901\}$, but the estimated probabilities were not markedly different for these other choices of $V$. With the exception of the IAC model, in this section we consider probabilities only for the complete ballot case.
5.1. Descriptions of the Models
We use two different types of models: one-dimensional spatial models and models which sample over the ballot simplex. In the spatial framework, we consider three different models. The first model, which we denote 1D Spatial, places three candidates and $V=1001$ voters on the real line using the standard normal distribution, and a voter’s preference ranking is determined by their Euclidean distance from each candidate. For example, if the three candidates $A$, $B$, and $C$ are positioned by the normal distribution at $-1.2$, $-0.1$, and 1.1 respectively, then a voter positioned at 0.3 would cast a ballot with $B$ ranked first, $C$ ranked second, and $A$ ranked third.
Because of the increased political polarization in the United States and elsewhere, we also use two bimodal spatial models to investigate the potential effects of polarization on the probability of the reinforcement paradox. Under the second spatial model, which we denote 1D Bimodal, we create a bimodal distribution from an equally weighted combination of the normal distribution centered at $-2$ and the normal distribution centered at 2, both with standard deviation 1. See Figure 1(top) for a visualization of this bimodal distribution. We then place candidates and voters along the line using the generated (normalized) bimodal distribution as a probability distribution, similar to 1D Spatial. Under the third spatial model, which we denote 1D Bimodal Weighted, we use the same two normal distributions to generate the bimodal distribution but give 50% more weight to the left distribution, resulting in a bimodal distribution as shown Figure 1(bottom). We use this model to capture polarization when one side of the political spectrum is stronger than the other.
Figure 1. Visualizations of the two bimodal spatial models.
For the ballot simplex models we use Dirichlet distributions with parameter $\alpha$, denoted Dir$(\alpha)$, and we use the implementation of these models provided by the software package VoteKit [10]. In a ballot simplex model, we place a probability distribution on the discrete ballot 5-simplex, which is the set of 6-tuples $(X_{1},X_{2},Y_{1},Y_{2},Z_{1},Z_{2})$ such that each number is non-negative and the six numbers sum to $V$. Under a Dirichlet model with $\alpha\in(0,\infty)$, as $\alpha\rightarrow 0$ the generated profile tends to weigh heavily in favor of a single candidate, so that the resulting election is a landslide victory for one of the candidates. For example, when $\alpha$ is close to zero, if $V=1001$ then a 6-tuple $(X_{1},X_{2},Y_{1},Y_{2},Z_{1},Z_{2})$ where $X_{1}+X_{2}>800$ is much more likely to be generated than a 6-tuple where each of the six numbers is approximately equal. When $\alpha=1$ the model places a uniform distribution on the simplex, so that any 6-tuple of ranking numbers is as likely as any other. Dir$(1)$ is referred to as the IAC model in social choice. As $\alpha$ increases the sampled 6-tuples tend to cluster in the center of the simplex, so that it becomes increasingly unlikely that a candidate earns a majority of first-place votes. In the limit $\alpha\rightarrow\infty$, the model is very similar to what is called the impartial culture (IC) model. Under the IC model, each voter independently chooses one of the six preference rankings at random, leading to elections that, with high probability, are closely tied. Thus, as $\alpha$ increases a typical generated election becomes highly competitive between the three candidates.
If we fix an $\alpha$, we can consider the limiting probability of observing a paradox where we let $V\rightarrow\infty$. In this limit, the discrete ballot simplex becomes the continuous 5-simplex determined by $X_{1}+X_{2}+Y_{1}+Y_{2}+Z_{1}+Z_{2}=1$, where each term in the sum is represents a (non-negative) proportion of votes. By using large electorates in our simulations, we estimate limiting probabilities under each model.
5.2. Results
Throughout this section and the next, in our analysis we identify four probabilities: the probability that a runner-up (RU) paradox occurs; the probability that a plurality loser (PL) paradox occurs; the probability that at least one of these (RU or PL) paradoxes occurs; and the probability that both (RU and PL) paradoxes occur.
We begin by providing exact probabilities under Dir$(1)$, the IAC model. This model is important in the social choice literature and therefore many researchers have dedicated significant effort to developing tools for calculating probabilities under this model. One of the fruits of these labors is the software package Normaliz [5], which can provide limiting probabilities under IAC in three-candidate elections given a set of inequalities (such as those given in Propositions 1 and 2). Normaliz can be used in the complete ballot case, where profiles $(X_{1},X_{2},Y_{1},Y_{2},Z_{1},Z_{2})$ are chosen at random from the 5-simplex, and in the partial ballot case, where profiles $(X_{1},X_{2},X_{3},Y_{1},Y_{2},Y_{3},Z_{1},Z_{2},Z_{3})$ are chosen at random from the 8-simplex. The probabilities in the following proposition were obtained directly from Normaliz by imputting the appropriate sets of inequalities.
Proposition 5.
The limiting probabilities of a reinforcement paradox occurring under the Dir$(1)$ model (the IAC model) are as shown in Table LABEL:IAC_results for complete and partial ballots. The “Cond. Prob.” probabilities are conditioned on the absence of a majority candidate.
PL or RU RU PL PL and RU
Prob. $\frac{191}{768}\approx 24.9\%$ $\frac{127}{576}\approx 22.5\%$ $\frac{79}{1152}\approx 6.9\%$ $\frac{31}{768}\approx 4.0\%$
Cond. Prob. $\frac{191}{336}\approx 56.8\%$ $\frac{127}{252}\approx 47.2\%$ $\frac{79}{504}\approx 15.7\%$ $\frac{31}{336}\approx 9.2\%$
Prob. $\frac{1341257}{5308416}\approx 25.3\%$, $\frac{527}{2304}\approx 22.9\%$ $\frac{7031}{110592}\approx 6.4\%$ $\frac{210439}{5308416}\approx 4.0\%$
Cond. Prob. $\frac{1341257}{3006720}\approx 44.6\%$ $\frac{527}{1305}\approx 40.4\%$ $\frac{7031}{62640}\approx 11.2\%$ $\frac{210439}{3006720}\approx 7.0\%$
Table 11. Probabilities under the Dir$(1)$ (IAC) model. The top two rows correspond to the complete ballot case; the bottom two rows correspond to the partial ballot case.
The probabilities of one or more paradoxes occurring are similar in the complete and partial ballots cases. When conditioned on the absence of a majority candidate, the probabilities are significantly lower in the partial ballot case. This is because the probability an election does not contain a majority candidate is $145/256=56.64\%$ in the 8-simplex but is only $7/16=43.75\%$ in the 5-simplex.
For finite $\alpha\neq 1$ we are unaware of analytical techniques to calculate exact probabilities, and therefore we use Monte Carlo simulation. Figure 2 illustrates the probabilities obtained from simulations for Dir$(\alpha)$ where $\alpha\in\{0.1,0.2,0.3,\dots,4.0\}$: the top image indicates unconditioned probabilities and the bottom image indicates probabilities conditioned on the absence of a majority candidate. The blue squares give the probability that an election demonstrates a reinforcement paradox (of either type), the red disks give the probability an election demonstrates a runner-up paradox, the black diamonds give the probability an election demonstrates a plurality loser paradox, and the green pluses give the probability an election demonstrates a double paradox.
Figure 2. Estimated probabilities that an election demonstrates a reinforcement paradox under the model Dir$(\alpha)$. The top figure shows unconditioned probabilities while the bottom figure shows probabilities conditioned on the absence of a majority candidate.
As $\alpha$ increases, all four probabilities also increase; this is unsurprising because as $\alpha$ increases, a typical generated election becomes “closer” or “more competitive.” When $\alpha$ is small, many generated elections represent landslide victories where one candidate has a strong majority of first-place votes. When $\alpha$ is large, the generated elections tend to cluster in the center of the simplex where there is no majority candidate. This also explains why we see a large increase in probabilities when moving from the top of Figure 2 to the bottom when $\alpha$ is small (when there are few elections without a majority candidate) but not when $\alpha$ is large (when almost all elections have no majority).
Table 12 provides specific results from the simulations for a few choices of $\alpha$. Recall that 100,000 simulations were run for each choice of $\alpha$, so that when $\alpha=0.1$, for instance, only 3.5% of generated elections contained no majority candidate. Note that we include simulations for $\alpha=1$ even though we have exact probabilities from Proposition 5.
Table 12 also displays information for three large $\alpha$ values which are not shown in Figure 2. Recall that as $\alpha\rightarrow\infty$ the distribution resembles the IC model, in which each voter independently chooses one of the six possible rankings at random. For comparison, we also conducted simulations using the IC model with $V=1001$, which returned an estimated probability of 1.0 for the runner-up, plurality, or both (RU, PL, and RU and PL) paradoxes occurring. Intuitively, this is clear since when every voter chooses a ranking at random, $X_{1}\approx X_{2}\approx Y_{1}\approx Y_{2}\approx Z_{1}\approx Z_{2}$. Thus, in the limit, Condition (2) and the plurality loser inequalities of Proposition 1 will always be satisfied. It is unsurprising, therefore, that the probabilities in Figure 2 appear to be trending toward 1 and the values at the bottom of Table 12 indicate convergence to 1 in all cases.
$\alpha$ PL or RU RU PL PL and RU No Maj. Cand.
0.1 1257 1033 242 18 3500
0.5 12419 10737 2730 1048 24024
1.0 24745 21852 6825 3932 43590
1.5 34319 31065 11082 7828 56455
2.0 42038 38548 15426 11936 65795
2.5 48079 44538 19721 16180 72663
3.0 53702 50123 23732 20153 78324
10.0 85618 83538 62924 60844 98498
20.0 96383 95392 86120 85219 99948
30.0 98814 98442 94367 93955 99997
$\infty$ (IC) 100000 100000 100000 100000 100000
Table 12. Simulation results for some choices of $\alpha$. For each $\alpha$ we generated 100,000 profiles.
The 1D spatial models provide another way to estimate the probabilities of a reinforcement paradox, in part because the preferences they represent are single-peaked. Thus, in simulations in which the positions of the candidates $A$, $B$ and $C$ are ordered left to right, for instance, $X_{2}=Z_{1}=0$ since $B$ must be either first or second-ranked by all voters.
The results under all spatial models are shown in Table 13. The first row gives unconditioned probabilities; the second row gives probabilities conditioned on the absence of a majority candidate for 1D spatial. The middle (respectively last) two rows give the corresponding probabilities under 1D Bimodal (respectively 1D Bimodal Weighted).
Interestingly, the corresponding probabilities under 1D Spatial and 1D Bimodal are approximately equal, perhaps because both models are symmetric about zero. When we unbalance the distribution under 1D Bimodal Weighted, the probabilities change significantly. In particular, 1D Bimodal Weighted is the only model in this study under which it is more likely to observe a plurality loser paradox than a runner-up paradox. We do not know with certainty the reason for this, but we speculate that it is due to different responses to the so-called “center squeeze” [39]. In one-dimensional three-candidate spatial models, the candidate lying in the center will often be the plurality loser even if they are closest to the “median voter.” Using the notation of Section 3, if $C$ is the center candidate then $X_{1}=Y+1=0$ since these voters will second-rank the nedian candidate. Hence the conditions in Case (1) of Proposition (1) cannot be satisfied (since they imply $B$ has the majority of the first-place vote). The conditions in Case (2) reduce to
$Y>X-Z_{2}\quad\mbox{ and }\quad Y>X_{2}+Z_{1}-Z_{2}=X+Z_{1}-Z_{2}.$
If the distribution of voters is symmetric, $Z_{1}$ and $Z_{2}$ are likely to be similarly-sized, and so the rightmost inequality will be satisfied whenever $Y>X$. Thus a runner-up (RU) paradox will occur often when the IRV and plurality winners are distinct.
If the distribution of voters is not symmetric, as with the 1D Bimodal Weighted model, the left and right-wing candidates are likely to be near the left and right peaks of the distribution respectively, with the centrist in between but on the “upward slope” of the larger peak (the left one in Figure 1). Thus, the quantity $Z_{1}-Z_{2}$ is more likely to be larger than in the symmetric case, decreasing the likelihood of Case (2) occurring.
A similar argument holds if $A$ or $B$ is the centrist candidate.
For the plurality loser (PL) paradox to occur, it is easily seen that $C$ must be the centrist candidate. In this case, $C$ is again likely to be on the upward slope of the larger peak in the 1D Bimodal Weighted model (and hence $Z$ is likely to be larger) than in the 1D Bimdodal model. So the probability of the plurality loser (PL) paradox occurring is greater in the nonsymmetric case than the symmetric.
Under all three 1D spatial models, the probability of a double paradox (both RU and PL) is extremely small as it is rare for both $Y$ and $Z$ to be sufficiently large simultaneously.
PL or RU RU PL PL and RU
Unconditioned Probability 16.63 12.17 4.47 0.01
Given No Majority Cand. 40.07 29.32 10.76 0.01
Unconditioned Probability 16.68 12.52 4.18 0.02
Given No Majority Cand. 34.05 25.55 8.53 0.03
Unconditioned Probability 11.46 5.05 6.46 0.05
Given No Majority Cand. 26.48 11.67 14.92 0.12
Table 13. Estimated probabilities, reported as percentages, under our three spatial models. The top two rows give probabilities for 1D Spatial, the middle two rows for 1D Bimodal, and the last two for 1D Bimodal Weighted.
6. Empirical Results
In this section we provide empirical results using a large database of real-world IRV elections. Our data comes from three sources. First, we use single-winner political elections from the United States with at least three (not write-in) candidates. This data is available at the FairVote data repository [34], and contains ballot data from municipal elections in cities such as San Francisco, CA and Minneapolis, MN. The repository also has ballot data for federal elections in the states of Alaska and Maine. Second, we use single-winner elections from the American Psychological Association (APA), which generally is willing to share ballot data for ranked-choice APA presidential elections and elections for the Board of Directors. Some of this data is available on preflib.org [23], and some was shared directly with the first author. Third, we use single-winner political elections from a database of Scottish local government elections. The ballot data for these elections is available at https://github.com/mggg/scot-elex.
In total, we accessed the preference profiles for 352 single-winner IRV elections with at least three candidates. [^4] Any reference to the number of candidates in an election does not include write-ins. In our dataset, write-in candidates receive an insignificant amount of votes and ignoring them does not affect the analysis. Many of these elections contain more than three candidates; because we are interested in the three-candidate case, for each election we “reduce” the number of candidates by running the IRV algorithm until only three candidates remain and then analyze the resulting profile. Note that the relationship between the original election and the resulting three-candidate election is not clear with respect to the susceptibility to reinforcement paradoxes. It is a priori possible that the original election demonstrates a paradox while the reduced three-candidate election does not, and vice versa. We use this methodology because we are interested in three-candidate elections and analyzing the elections with more than three candidates in this way the IRV algorithm in this allows us to expand the size of our three-candidate election dataset. Moreover, in some cases if the resulting three-candidate election demonstrates a paradox then we can gain insight into how the original election demonstrates a paradox; this is how a reinforcement paradox in a five-candidate Minneapolis city council election was found, for example [25].
Jurisdiction USA APA Scotland Total
Elections 295 27 30 352
3-Cand. Elections 104 10 1 115
3-Cand. Elections, no Majority Cand. 44 6 1 51
3-Cand. Reduced Elections, no Maj. Cand. 164 20 24 208
Table 14. A summary of our data sources.
6.1. Results
Table 15 provides the results for elections having three candidates in the original election and for all 352 elections. We use the same abbreviations RU and PL as in Section 5. The first row indicates the number and percentage of elections exhibiting one or more paradoxes among the three-candidate elections. For example, of the 115 elections which originally contained three candidates, 17 of them (14.8%) exhibit a reinforcement paradox. The second row indicates the same information but for all the elections including the reduced elections. The third and fourth rows indicate the percentage of elections that exhibit the given type of paradox, conditioned on the absence of a majority candidate.
Unlike the simulated elections in Section 5, all the elections in the database that demonstrate a plurality loser paradox also demonstrates a runner-up paradox. (Note that the two rightmost columns of Table 15 are identical.) This is because in real elections runner-up candidates tend to be much stronger than plurality loser candidates.
PL or RU RU PL PL and RU
3-Cand. Elections 17, 14.8% 17, 14.8% 7, 6.1% 7, 6.1%
3-Cand. Reduced Elections 86, 24.4% 86, 24.4% 27, 7.7% 27, 7.7%
3-Cand. Elections, no Maj. Cand. 33.3% 33.3% 13.7% 13.7%
3-Cand Reduced Elections, no Maj. Cand. 41.3% 41.3% 13.0% 13.0%
Table 15. Reinforcement paradox results for the 115 real-world three-candidate elections and for all 352 “reduced” elections.
In Section 4, we showed in Proposition 4 that if an election demonstrates a runner-up reinforcement paradox then the runner-up could receive as little as 25% of the first-place votes and the plurality loser could receive an arbitrarily small percentage of the first-place votes. Similarly, if an election demonstrates a plurality loser paradox then the plurality loser can receive as little as 25% of the first-place votes. How close to these minima do we come in real-world elections? The lowest percentage of first-place support earned by the plurality loser in an election which demonstrated a reinforcement paradox is 3.6%, which occurred in the 2023 city council ward 7 election in Minneapolis. Table 16 shows the preference profile for this election, which originally contained three (not write-in) candidates; that is, the profile displayed is not obtained from eliminating previous official candidates. The three candidates were Cashman ($C$), Graham ($G$), and Foxworth ($F$). Foxworth earns only 280 (3.6%) first-place votes of the 7969 ballots cast, yet the runner-up inequalities in Condition (2) of Proposition 2 are satisfied.
Num. Voters 1440 1217 1212 1473 611 1726 185 68 37
1st $C$ $C$ $C$ $G$ $G$ $G$ $F$ $F$ $F$
2nd $G$ $F$ $C$ $F$ $C$ $G$
3rd $F$ $G$ $F$ $G$ $G$ $C$
Table 16. The preference profile for the 2023 city council ward 7 election in Minneapolis.
This 2023 Minneapolis city council election is an interesting example of a reinforcement paradox, even though it initially appears unremarkable from an IRV perspective. In this election the plurality, Condorcet, and IRV winners are identical, ruling out the possibility of classic issues like a monotonicity paradox, spoiler effect, or similar pathology. While the election is close—Cashman and Graham receive a comparable number of first-place votes—Cashman is the clear winner. From a social choice perspective, this election typically warrants no further analysis. However, the presence of a reinforcement paradox makes this election more compelling than it might appear at first glance.
The lowest first-place vote percentage earned by the IRV runner-up in an election which demonstrates a runner-up paradox is 28.9%, achieved in the 2018 mayoral election in San Francisco. This election had 9 official candidates, and thus the 28.9% represents the vote share of the runner-up after eliminating six candidates. The lowest percentage for a plurality loser in an election which demonstrates a plurality loser election is 26.9%, achieved in the 2008 County Executive election in Pierce County, Washington. This election had four official candidates, and the plurality loser earned 26.9% of the voter after eliminating one candidate.
6.2. Bootstrapping Results
In Section 5 we ran simulations based on sampling using theoretical models such as the IAC model. Since we have access to a large dataset of real-world elections, we can also sample from these elections and run simulations using nonparametric bootstrap analysis. To perform this analysis, we follow techniques used in the classical papers [38] and [40]. For each of the 352 three-candidate profiles we have from our real-world dataset, we created 10,000 pseudoprofiles by sampling ballots with replacement from the actual three-candidate profile. This sampling allowed us to generate 3,520,000 three-candidate elections which we can check for reinforcement paradoxes. To keep the computation time reasonable and to match the analysis in Section 5, if the actual profile contained more than 1,001 voters then for each generated pseudoprofile we sampled only 1,001 ballots; otherwise each pseudoprofile contained the same number of voters as the original election. Most of the elections in the dataset contain more than 1,001 voters and thus most of the generated elections contain 1,001 voters.
Table 17 shows the probability of each type of reinforcement paradox across the 3,520,000 pseudoprofiles. The results are similar to the “3-Cand Reduced Elections” probabilities in Table 15, and thus the bootstrapping results do not meaningfully affect how frequently we expect such paradoxes to occur in real-world elections.
However, bootstrapping analysis can give a sense of how “close” a given election comes to demonstrating a paradox. If none of the 10,000 pseudoprofiles generated from an election demonstrate a paradox then the original election is not “close” to an election which demonstrates a paradox. On the other hand, if an election demonstrates a paradox but relatively few of its pseudoprofiles do, then the election is close to one which does not demonstrate a paradox. For example, the 2023 city council ward 7 election in Minneapolis (Table 16) demonstrated a reinforcement paradox but only 3370 of 10000 generated pseudoprofiles demonstrate a paradox. Thus, we could argue that this election demonstrating a paradox is not the most likely outcome.
A more nuanced summary of the bootstrapping results can be seen in Figure 3. Each histogram illustrates the frequency of the proportion of pseudoprofiles generated by each original election that resulted in a “paradoxical” election. The top image is based on the 86 original three-candidate elections that demonsrated the paradox. Thus, for example, the right-most bar on the top histogram indicates that in 33 of these elections, between 95 and 100% of the associated pseudoprofiles exhibited the paradox. Similarly, the second-to-the-right bar indicates that in 10 of these elections, between 90 and 95% of the associated pseudoprofiles exhibited the paradox. The bottom histogram is constructed the same way for the 266 original three-candidate elections which did not exhibit the paradox with the exception that elections which generated zero “paradoxical” pseudoprofiles are omitted for clarity (since the left bar would be out-sized).
From Figure 3 (top), we can observe that roughly half of the original elections exhibiting a paradox were extremely likely to exhibit that paradox in the sense that more than 90% of their associated pseudoprofiles also did so. Most of the remaining elections in this category also generated pseudoprofiles that more often than not exhibited a reinforcement paradox. The reverse is true in Figure 3 (bottom): the large majority of the original elections that did not exhibit the paradox generated few pseudoprofiles exhibiting a reinforecement paradox.
Noteworthy, however, is the fact that across all 352 original elections, 206 generated at least one pseudoprofile which demonstrated a reinforcement paradox. Thus, arguably, 206 elections are close to an election which exhibits a paradox. If we restrict the analysis to elections that did not contain a majority candidate, 187 of the 208 of the original elections generated at least one pseudoprofile exhibiting a reinforcement paradox. These results suggest that almost any three-candidate real-world election which does not contain a majority candidate is “close” to an election demonstrating a reinforcement paradox. We note such a high percentage is not achieved under bootstrapping for other paradoxes. For example, in [27] the same type of analysis was performed for the spoiler effect using a smaller dataset of American IRV elections (but still using 10,000 pseudoprofiles per election), and only five elections generated at least one pseudoprofile demonstrating a spoiler effect under IRV. Thus, it is not the case that performing this bootstrap analysis guarantees that almost all elections in a dataset will be “close” to an election which exhibits a given paradox.
Recall in the 352 actual elections, any election demonstrating a plurality loser paradox also demonstrated a runner-up paradox. This is not the case for pseudoprofiles: 9,542 generated elections demonstrate a plurality loser but not a runner-up paradox. Thus, while such an outcome is not likely, some real-world elections are close to this phenomenon occurring.
PL or RU RU PL PL and RU
Unconditioned Prob. 26.22% 23.94% 7.21% 6.94%
Conditioned Prob. 42.29% 41.82% 12.60% 12.13%
Table 17. Probability, expressed as a percentage, of observing a reinforcement paradox in the 3,520,000 generated pseudoprofiles.
Figure 3. (Top) Bootstrapping frequency results for elections which demonstrated a reinforcement paradox. (Bottom) Bootstrapping frequency results for elections which did not demonstrate a reinforcement paradox but generated at least one pseudoprofile that did.
7. Discussion
In this section we discuss our results, starting with a comparison between the theoretical and empirical results, then placing our results into context by comparing them to other classical paradoxes, and concluding by comparing our results to previous results on the reinforcement paradox.
7.1. Theoretical Versus Empirical Results
As is often the case in social choice, none of the theoretical models give probabilities which closely match what we observe in real-world elections across all four paradox types. Interestingly, if we focus on the 352 three-candidate reduced elections (lines two and four of Table 15), the IAC complete-ballot model (Table LABEL:IAC_results) provides a decent approximation of the empirical results. Furthermore, even though the real-world elections contain many partial ballots, generally the complete-ballot IAC model more accurately reflects the empirical probabilities across the 352 reduced elections than the partial-ballot IAC model. For other voting paradoxes such as monotonicity paradoxes or the spoiler effect, the IAC model typically gives probabilities much higher than what we observe in real elections. If we focus only on the 151 three-candidate actual elections, Dir$(0.6)$ gives the best approximation of the empirical unconditioned results (with the exception of “PL and RU”), which makes sense because in these elections, the winner is stronger overall than in the Dir$(1)$ (IAC) model, consistent with the election data. Overall, the 1D-spatial models, both normal and bimodal, underestimate the probabilities of one or more types of paradox occurring.
What none of the models capture is that in actual elections, it is extremely rare to observe a plurality loser paradox without also observing a runner-up paradox. Under spatial models this difference is especially pronounced, as under those models if a plurality loser paradox is exhibited then we almost never also observe a runner-up paradox. This is generally true across all models, even though the runner-up (RU) paradox is roughly three times as likely to occur as the plurality loser (PL) paradox.
All theoretical models predict that if an election is somewhat “competitive” then the probability of demonstrating a reinforcement paradox is large (although usually not larger than 50%). The empirical results also bear this out, with empirical probabilities sometimes being larger than theoretical ones, even when using theoretical models which often produce close elections. Of course, the empirical probabilities come nowhere close to the probability of 1 predicted by IC, the model which tends to produce extremely competitive elections.
7.2. Comparison to Other Paradoxes
Compared to other classical paradoxes such as downward and upward monotonicity paradoxes, no-show paradoxes, or the spoiler effect, the probability an election demonstrates a reinforcement paradox is much larger. For example, Graham-Squire and McCune [17] use a very similar database to provide empirical probabilities, conditioned on the absence of a majority candidate, of 2.2%, 1.6%, 0.5%, and 2.2% for upward monotonicity paradoxes, downward monotonicity paradoxes, no-show paradoxes, and the spoiler effect, respectively. The unconditioned probabilities fall to around 1% or less. Theoretical probabilities for these other paradoxes under models like IAC or IC also tend to be much lower than what we observe for reinforcement paradoxes [18, 27]. In particular, we are not aware of any other paradox for which the IC model returns a probability of 1.
As far as we are aware, prior to this work the paradox which gave the highest probabilities is the negative participation paradox, explored in [18, 28]. McCune and Wilson [28] give an unconditioned empirical probability of 3.3%, which rises to 4.9% when we condition on the absence of a majority candidate. They also show that for a very large electorate with complete ballots, the unconditioned probability of observing this paradox is 15.77% under the IC model, and Kamwa et al. [18] give a probability of 7.69% under the IAC model. These probabilities are significantly smaller than the reinforcement paradox probabilities.
7.3. Comparison to Previous Reinforcement Paradox Results
As mentioned in the Introduction, we are aware of only two previous papers [8, 37] which estimate the probability with which an election demonstrates a reinforcement paradox. Our work differs from theirs in several significant ways. Most importantly, in most cases, the analysis in these earlier papers was based on simulating two separate elections, checking whether the elections had the same winner, and then indicating a reinforcement paradox had occurred if the winner obtained from the combined election was different from the winner in the two generated elections. That is, the authors did not start with one election and look for a partition of the ballots as we do here. As Plassmann and Tideman state in [37], “We determine the frequencies of the reinforcement paradox by sampling ballots for two separate elections, which we then combine to determine whether the paradox occurred for any voting rule. In contrast, we determine the frequencies of all other paradoxes from a single sample of ballots.” Plassmann and Tideman perhaps made this choice from convenience, as they later state, “There is no obvious way in which to divide the ballots from a single election into two distinct sets that would yield the appropriate distribution of rankings in the two new sets and would permit us to examine reinforcement.” We have contributed to their work by providing such a way to divide the ballots in the 3-candidate case in Propositions 1 and 2. Courtin et al. [8] generally take a similar approach, simulating two ballot sets and then combining them. We feel our approach is equally valid since we are interested in the “paradoxical harm” felt by a candidate (and perhaps their supporters) in an election exhibiting the paradox: if there exists a partition of the electorate such that one of the losing candidates can point to each sub-electorate and claim victory, that candidate can reasonably claim some sort of harm when they ultimately lose the overall election.
A second way our work differs from previous studies is that we use a much larger set of theoretical models as well as empirical data to conduct our analysis. Neither of the previous papers uses any real-world ranked-choice ballot data. [^5] The paper by Courtin et al. does have a section where they apply their results to parliamentary (or House) elections in France, Germany, and the United States, but the authors do not have access to candidate preference rankings. Our use of multiple spatial models, multiple Dirichlet models, and a large dataset of real-world elections (combined with nonparametric bootstrapping analysis) greatly expands previous work.
Thirdly, ours is the first work to provide any analysis of partial ballots in the context of reinforcement paradoxes. Previous work was not concerned with partial ballots, presumably because such work did not analyze any real-world IRV elections.
Finally, both previous papers report extremely small probabilities for observing a reinforcement paradox under IRV. Plassmann and Tideman give a probability of approximately 0.02% when using an electorate of at least 1,000 voters. Courtin et al. report similarly small probabilities. Using the term “Young’s reinforcement axiom” (YRA) synonymously with Young’s notion of consistency (Definition 1), Courtin et al. state, “Our results show that … cases of violations of YRA are rather rare… [[[[A]]]]lthough all these [[[[iterative scoring]]]] rules violate YRA, the frequencies of violation are very rare. Then, since the violation of this axiom is not so frequent, one should not be unduly worried about its theoretical possibility.” By contrast, we report extremely large reinforcement paradox probabilities, which are generally much larger than the probabilities of observing other classical paradoxes. The reason for this discrepancy is our difference in approach, as outlined above. We would also obtain similarly small probabilities if we were to take each real election and search for reinforcement paradoxes by choosing a partition of the ballots at random, for example. We obtain higher probabilities because we control how the ballots are partitioned.
8. Conclusion
Both theoretical and empirical analysis suggests that the reinforcement paradox is significantly more likely to occur under IRV with 3 candidates than other paradoxes. This is partly because, as noted above, there are no restrictions on how the ballots are partitioned. If we require the two sub-elections to be similarly-sized, for example, as they might be in two precincts, there will most likely be fewer instances of the paradox; the same is true if we restrict the sub-elections so that one sub-election is a small perturbation of the original election. Future work could examine how the probability of a reinforcement paradox changes under restrictions of this kind, or to instances involving a larger number of sub-elections. Finally, this work could be extended to more than three candidates by determining conditions on the parameters in an $n$-candidate preference profile under which the paradox occurs.
The results in this paper also shed light on ways in which IRV differs from positional scoring methods, which are not susceptible to the reinforcement paradox. It would be interesting to repeat this analysis for other voting methods, such as Condorcet consistent methods, that do not fall into this class.