### v3

# Instant Runoff Voting and the Reinforcement Paradox

Abstract.
We analyze the susceptibility of instant runoff voting (IRV) to a lesser-studied paradox known as a reinforcement paradox, which occurs when candidate $X$ wins under IRV in two distinct elections but $X$ loses in the combined election formed by merging the ballots from the two elections. For three-candidate IRV elections we provide necessary and sufficient conditions under which there exists a partition of the ballot set into two sets of ballots such that a given losing candidate wins each of the sub-elections. Applying these conditions, we use Monte Carlo simulations to estimate the frequency with which such partitions exist under various models of voter behavior. We also analyze the frequency with which the paradox occurs in a large dataset of real-world ranked-choice elections to provide empirical probabilities. Our general finding is that IRV is highly susceptible to this paradox in three-candidate elections.
Key words and phrases:
instant runoff voting, ranked choice voting, reinforcement paradox, consistency, empirical results
The voting method of instant runoff voting (IRV), also referred to as the alternative vote, the plurality elimination rule, and colloquially as “ranked-choice voting” in the United States, is famously susceptible to many classical voting paradoxes [22, 42, 52]. In this article we analyze the susceptibility of IRV to a lesser-studied paradox known as a reinforcement paradox, which occurs when candidate $X$ wins under IRV in two distinct elections but $X$ loses in the combined election formed by merging the ballots from the two elections. For three-candidate IRV elections we provide necessary and sufficient conditions under which there exists a partition of the ballot set into two subsets of ballots such that a given losing candidate wins each of the sub-elections. Applying these conditions, we use Monte Carlo simulations to estimate the frequency with which such partitions exist under various models of voter behavior. We then analyze the frequency of the paradox in a large dataset of real-world ranked-choice elections to provide empirical probabilities. By characterizing the three-candidate IRV elections that demonstrate susceptibility to the reinforcement paradox and estimating its frequency, we view our work as a contribution to the larger agenda laid out by Felsenthal in [20].
Reinforcement paradoxes have received attention in the social choice literature due to their interesting and surprising nature, much like other paradoxes to which IRV is susceptible, such as upward or downward monotonicity paradoxes. However, unlike monotonicity paradoxes, IRV’s susceptibility to reinforcement paradoxes has direct administrative implications: to determine the IRV winner, it is not sufficient to know the winner at each precinct. Instead, complete information about the numbers of ballots with each different ranking must be transmitted—either electronically or physically—to a central counting location and collated before the IRV winner can be determined. In contrast, a voting method such as plurality (or a scoring rule) allows each precinct to simply report the number of first-place (and possible second-place, etc.) votes for each candidate, and for the winner to be determined by summing these results across precincts. As a result, IRV ballot tabulation and the subsequent determination of the winner is of necessity more complicated than with methods like plurality. Thus, a study of reinforcement paradoxes is motivated by practical as well as theoretical considerations.
Our work in this paper connects to three previous strands of the literature concerning IRV and voting paradoxes. The first gives conditions on a preference profile which characterize when an election is susceptible to a given paradox. Examples are Lepelley et al. [31], Miller [39], and Ornstein and Norman [43], which characterize the three-candidate profiles which demonstrate an upward or downward monotonicity paradox under IRV. Similarly, Graham-Squire [24] characterizes three-candidate profiles which demonstrate a no-show paradox under IRV, including profiles which contain partial ballots. Kamwa et al. [27] study a variety of voting rules, but in particular provide conditions under which three-candidate profiles produce paradoxes such as negative participation paradoxes under IRV.
The second strand of literature typically uses the conditions given by the works cited above to estimate the frequency with which a given paradox occurs under IRV. For example, Ornstein and Norman [43] estimate the frequency of monotonicity failures in IRV elections when the ballot data is generated using a spatial model. Similarly, Lepelley et al. [31] and Miller [39] estimate the frequency with which IRV produces monotonicity paradoxes under various theoretical models of voter behavior. Lepelley and Merlin [32] analyze the frequency with which IRV demonstrates a no-show paradox in three-candidate elections under the impartial culture and impartial anonymous culture models. Kamwa et al. [27] analyze the probability that IRV (among other methods) demonstrates no-show or negative participation paradoxes on single-peaked domains. McCune and Wilson [38] examine the susceptibility of IRV to negative participation paradoxes.
More broadly, there is a large literature examining the extent to which various voting rules are susceptible to strategic manipulation. Such work need not rely on algebraic characterizations involving the vote totals in the preference profile; see [1, 19, 26, 28] for such examples.
The third strand is the expanding empirical social choice literature on IRV, enabled by the increasing availability of ranked-choice ballot data from jurisdictions such as Australia [54], the United States [44], and Scotland [34]. Some of this literature concerns the empirical probability that IRV produces a particular kind of paradox. Graham-Squire and Zayatz [23], Graham-Squire and McCune [25], and McCune and Graham-Squire [34] estimate the frequency with which real-world IRV elections demonstrate monotonicity and no-show paradoxes. Other empirical work includes case studies of particularly interesting elections which demonstrate a range of different paradoxes; see [30, 35], for example. Beyond work on paradoxes, empirical research on IRV has analyzed rates of ballot errors in IRV elections [11, 45], the extent of ballot exhaustion [10, 25], the rate at which IRV exhibits the so-called “spoiler effect” [37], the extent to which ballot truncation can affect the IRV winner [14, 29, 56], and the frequency with which IRV selects the Condorcet winner [25, 53, 54]. There is also a growing, related empirical literature concerning the method of the single transferable vote (STV), a multiwinner version of IRV which is designed to achieve proportional representation given a set of preference ballots. This literature also investigates paradoxical and Condorcet-related outcomes [34, 36], as well as STV’s propensity to achieve proportional representation [4, 6, 40].
Of course, there is a much broader literature concerning paradoxes and voting methods other than IRV, and we do not attempt to survey that literature here. See [16, 41, 46] for an examination of the susceptibility of Condorcet consistent methods to various types of no-show paradox, for example.
Previous work on the reinforcement paradox is fairly limited. The notion of a reinforcement paradox dates to the work of Young [57, 58], who defined the concept and proved essentially that the only reasonable voting methods not susceptible to the paradox are positional scoring rules. Felsenthal and Nurmi [21] show that various Condorcet inconsistent voting rules (including IRV) are susceptible to the paradox even in “tame” profiles in which a Condorcet winner exists and the voting rule chooses this candidate. Brandt et al. [5] provide some “minimal” examples of the reinforcement paradox under IRV where the number of candidates and voters is small. Brandt et al. [8] analyze the frequency of reinforcement paradoxes under various Condorcet consistent voting rules. We are aware of only two previous papers which attempt to measure the frequency with which IRV produces reinforcement paradoxes, both of which focus on the three-candidate case. Plassmann and Tideman [47] measure this frequency under a spatial model, and Courtin et al. [12] study the paradox under the impartial culture and impartial anonymous culture models of voter behavior. Our work differs from these studies in significant aspects which we explicate more fully below. In brief, the approach taken by these studies is to generate two elections and then say that the reinforcement paradox has occurred if the two elections have the same winner, but when the ballots from the two elections are combined, the winner is different. By contrast, we consider only one election and say that the paradox has occurred in this election if there exists a way to partition the ballots into two sub-elections where the winners of the sub-elections are the same but differ from the winner of the original election. Our work is complementary to theirs and allows us to focus on one set of ballots and the characteristics required for it to demonstrate a susceptibility to this kind of inconsistency.
Another reason for our interest in this approach is that it parallels a well-established line of research regarding self-consistency in allocation rules. For instance, Aumann and Maschler [2] identify consistency (in which the outcome of dividing assets among a set of agents doesn’t change when applied to a subset of agents) as a crucial characteristic of claims methods. Balinski and Young [3] define an analogous notion of uniformity to characterize divisor methods of apportionment. Converse consistency and other related ideas also appear in the cooperative game and bargaining literatures; see [55] for a survey.
Consistency in allocation rules is inherently different from consistency in voting rules: consistency in allocation rules requires that if an allocation is recalculated after a subset of agents “departs” with their share, the allocation does not change. Put another way, if the agents are partitioned into two subsets and the shared goods in each subset are pooled and then reallocated, the results are unchanged. We can adapt this latter interpretation loosely to the social choice framework: if an election is partitioned into two subsets and the winner is determined in each of the subsets, the results should reflect the original winner.
The paper is structured as follows. Section 1 gives preliminaries, including formal definitions of IRV and the reinforcement paradox. In Section 2 we provide necessary and sufficient conditions for a three-candidate election to demonstrate a reinforcement paradox and discuss some of the consequences of these conditions. Proposition 1 gives these conditions assuming all voters provide a complete ranking of the candidates while Proposition 2 gives conditions when some voters cast partial ballots. Section 3 considers the minimum percentage a losing candidate can earn and yet the election demonstrates a reinforcement paradox. Section 4 estimates the probability that an election demonstrates a paradox under various theoretical models of voter behavior. Section 5 analyzes real-world IRV elections from the United State and Scotland. This section provides empirical probabilities using a large dataset of elections, and provides a probabilistic analysis using non-parametric bootstrapping. Section 6 gives a discussion of our results, and Section 7 concludes.
1. Preliminaries
In this section we define IRV and different types of reinforcement paradoxes. We illustrate these ideas with several examples.
In an election which uses IRV, voters cast preference ballots with a (possibly partial) linear ranking of the candidates. After an election, the ballots are aggregated into a preference profile, which shows the number of each type of ballot cast. For example, Table 1 shows a preference profile involving the three candidates $C$, $G$, and $H$. The number 2211 denotes that 2211 voters ranked $G$ first, $H$ second, and $C$ third. In the real-world IRV elections we analyze in Section 5, voters are allowed to provide partial rankings. Table 1 has three columns where voters do not provide a complete ranking–2003 voters rank only $G$ on their ballots, for example.
An IRV election proceeds in rounds where, in each round, the number of first-place votes for each candidate is calculated. If a candidate has a majority of the (remaining) first-place votes, they are declared the winner. If no candidate receives a majority of votes, the candidate with the fewest first-place votes is eliminated and the names of the remaining candidates on the affected ballots are moved up. The process repeats until a candidate is declared the winner. In the case when partial ballots are allowed, any ballot in which all candidates have been eliminated is considered “exhausted” and plays no further role in the determination of the winner. [^1] In theory there are many ways we could treat exhausted ballots besides removing them from the election. However, as far as we are aware, any actual election office treats exhausted ballots in this fashion.
When there are only three candidates, at most two rounds are required to determine a winner; in general, several rounds may be required. We focus only on the three-candidate case and thus do not consider more than two rounds of ballot counting. The following example demonstrates the IRV algorithm using the 2022 Alaska Senate District E election in Alaska.
Num. Voters 2211 1439 2003 2605 241 2689 2230 418 2873
1st choice $G$ $G$ $G$ $H$ $H$ $H$ $C$ $C$ $C$
2nd choice $H$ $C$ $G$ $C$ $G$ $H$
3rd choice $C$ $H$ $C$ $G$ $H$ $G$
Table 1. Vote totals in the 2022 Senate District E election in Alaska.
Example 1.
(2022 Alaska Senate District E.) This election contained the three candidates Roselynn Cacy (C), Cathy Giessel (G), and Roger Holland (H). The preference profile is displayed in Table 1.222Our vote totals differ from the officially reported vote totals by up to three votes. This minor discrepancy could be due to a recount which occurred after the election, or because the Alaska Division of Elections has not been completely clear about how it processes “edge case” ballots. Our analysis is unaffected.
Initially, the vote totals for Cacy, Giessel, and Holland are 5521, 5653, and 5535, respectively. Cacy has the fewest votes and is eliminated from the election. As a result, 2230 votes are transferred to Giessel and 418 to Holland (the 2873 ballots ranking only Cacy are transferred to no one), and Giessel wins in the last round 7883 votes to 5953.
While Giessel is the winner, this election has the strange property that Cacy can point to two disjoint groups of voters such that Cacy would win under IRV if we use only the ballots from each group. Similarly, Holland can also break the electorate into two disjoint groups such that Holland wins when using the ballots from each group separately. We illustrate this paradoxical behavior in the next example.
Example 2.
In the 2022 Senate District E election in Alaska, suppose we partition the ballots into two ballot sets $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ as shown for Partition 1 in Table 2. If we use just the ballots from $\mathcal{B}_{1}$, $H$ wins under IRV, as $C$ is eliminated first and enough votes are transferred to $H$ for him to win against $G$ win in the final round. Similarly, if we use only the ballots from $\mathcal{B}_{2}$ then $H$ also wins under IRV, as in this case $G$ is eliminated first and enough votes are transferred to $H$ for him to win against $C$ in the final round. Thus, Roger Holland is able to point to two disjoint sub-electorates in which he would win the election using just the ballots from each sub-electorate, yet when we merge the ballots from the sub-electorates to recover the original ballot data, Holland loses.
Table 2 provides a second partition of the ballots, Partition 2, in which $C$ wins under IRV when using only the ballots from $\mathcal{B}_{1}$ and only the ballots from $\mathcal{B}_{2}$, yet when we merge the ballots to recover the original election data, $C$ loses. For both Cacy and Holland, in this election win ++ win == loss in some sense.
Num. Voters 2211 1439 2003 2605 241 2689 2230 418 2873
1st choice $G$ $G$ $G$ $H$ $H$ $H$ $C$ $C$ $C$
2nd choice $H$ $C$ $G$ $C$ $G$ $H$
3rd choice $C$ $H$ $C$ $G$ $H$ $G$
Partition 1 $\mathcal{B}_{1}$ 0 1439 1500 0 0 2689 0 418 2200
$\mathcal{B}_{2}$ 2211 0 503 2605 241 0 2230 0 673
Partition 2 $\mathcal{B}_{1}$ 0 1439 0 2605 0 0 0 0 1440
$\mathcal{B}_{2}$ 2211 0 2003 0 241 2689 2230 418 1433
Table 2. Two different partitions of the ballots in the 2022 Senate District E election in Alaska.
Example 2 demonstrates the notion of a reinforcement paradox, the subject of this article. The study of this paradox dates to the work of Young [57]. Young defined a voting method as a function $f$ which takes as input a set of ranked ballots and which outputs an election winner (or a set of winners, if there is some sort of tie). We use a simplified version of Young’s definition which does not address the issue of tied outcomes because ties almost never arise in the elections we study. In particular, none of our real-world elections have a tied outcome in any round. If $\mathcal{B}$ is a set of ballots and $f(\mathcal{B})$ is the winner under voting method $f$, then Young defined the following notion of consistency [57].
Definition 1.
Let $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ be two sets of ballots over the same candidate set. A voting method $f$ is said to be consistent if whenever $f(\mathcal{B}_{1})=f(\mathcal{B}_{2})$ then $f(\mathcal{B}_{1}\cup\mathcal{B}_{2})=f(\mathcal{B}_{1})$.
A voting method that is not consistent is said to be susceptible to the reinforcement paradox. This paradox has other names, such as the “multiple districts paradox” [22], and is a social choice analogue of the classical Simpson’s Paradox from statistics. In [58], Young showed that consistency and a few other uncontroversial axioms completely characterize positional scoring rules. Thus, any “reasonable” voting method which is not a scoring rule is not consistent. We analyze only IRV and the function $f$ represents that voting method throughout the article.
To investigate the frequency with which reinforcement paradoxes occur, we must first decide what constitutes an “occurrence” or an “instance” of the paradox. There seem to be two viable approaches. The first, used by Courtin et al. [12] and Plassman and Tideman [47], focuses on the ballot sets $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$. Their work generates ballot sets $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ independently using some probability distribution, and the paradox is said to occur if the election winner using $\mathcal{B}_{1}$ is the election winner using $\mathcal{B}_{2}$, but we obtain a different winner for the set $\mathcal{B}_{1}\cup\mathcal{B}_{2}$. A second approach is to start with a ballot set $\mathcal{B}$ and determine if it is possible to decompose into two subsets $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ where the winners of $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ match but differ from the winner of the original election. That is, in the second approach we select a single ballot set $\mathcal{B}$ using some probability distribution (or drawing from a set of real-world elections), and say this single election is susceptible to the reinforcement paradox if $\mathcal{B}$ can be partitioned into two sub-elections $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ such that $f(\mathcal{B}_{1})=f(\mathcal{B}_{2})$ but that the winner of the two sub-elections is different from $f(\mathcal{B}).$
We do not think either approach is “better” than the other; we view them as complementary. We choose to adopt the second approach for the following reasons. First, we are motivated by a study of real-world elections (see Section 5), and given a single actual election the second approach seems natural. If we focus on decomposing one ballot set into a partition then we can meaningfully ask “Did the 2022 Alaska Special House Election demonstrate susceptibility to a reinforcement paradox?”, and the answer is Yes if we can find a suitable partition of the election’s ballot set.
Second, this approach is in line with notions of “occurrence” for other paradoxes such as monotonicity paradoxes. For example, informally an election is said to demonstrate an upward monotonicity paradox if there exists a set of voters who could turn the winning candidate into a loser by shifting the winner to higher rankings on their ballots [23, 25]. That is, given a single election we demonstrate this paradox by finding a subset of ballots which, when suitably changed, produces a paradoxical outcome. Our approach with reinforcement paradoxes has a similar flavor: while we do not change the ballots, we hunt for subsets of ballots which produce a paradoxical outcome.
Third, much of the interesting work for the first approach has already been covered in [12, 47], but the second approach has not been well-studied. Brandt et al. [5] give results about “minimal” reinforcement paradoxes using the second approach; for example, for a fixed number of candidates and for various voting methods, they give a profile with a minimal number of voters such that the preference profile can be partitioned into two ballots sets which demonstrate the paradox. But otherwise we are unaware of previous results about partitioning a single ballot set. There is much interesting mathematics to explore in the second approach, such as providing necessary and sufficient conditions for an election to admit an appropriate partition of the ballot set. (We do exactly this, for three candidates, in Section 2.) For these reasons, we adopt the following definition for an election to demonstrate a reinforcement paradox.
Definition 2.
An election with ballot set $\mathcal{B}$ is susceptible to a reinforcement paradox under a voting method $f$ if there exists a partition of the ballot set $\mathcal{B}=\mathcal{B}_{1}\cup\mathcal{B}_{2}$ such that $f(\mathcal{B}_{1})=f(\mathcal{B}_{2})$ and $f(\mathcal{B}_{1})\neq f(\mathcal{B}_{1}\cup\mathcal{B}_{2})$.
Under IRV, if an election is susceptible to a reinforcement paradox then the paradox’s dynamics are similar to that of Example 2, in the sense that the elimination order in the two sub-elections are different. It is straightforward to show that if the IRV election using ballots from $\mathcal{B}_{1}$ has the same candidate elimination order as the IRV election using ballots $\mathcal{B}_{2}$ then the winner when using $\mathcal{B}_{1}\cup\mathcal{B}_{2}$ is necessarily the winner of the two sub-elections. Thus, when searching for a partition to demonstrate the paradox with three candidates, we look for sub-elections in which the two candidates in the final round of the election using $\mathcal{B}_{1}$ are not the same as the two candidates in the final round of the election using $\mathcal{B}_{2}$.
When there are three candidates an election can demonstrate a reinforcement paradox in two different ways, leading to the two definitions below. We refer to the candidate with the smallest number of first-place votes as the plurality loser. This candidate can never be the winner in an IRV election since they are eliminated in the first round. We refer to the remaining losing candidate, (who may receive either the largest or second-largest number of first-place votes) as the runner-up.
Definition 3.
An election is susceptible to a plurality loser reinforcement paradox if there exists a partition of the ballots $\mathcal{B}=\mathcal{B}_{1}\cup\mathcal{B}_{2}$ such that the plurality loser is the IRV winner when using only the ballots in $\mathcal{B}_{1}$ and when using only the ballots in $\mathcal{B}_{2}$.
Definition 4.
An election is susceptible to a runner-up reinforcement paradox if there exists a partition of the ballots $\mathcal{B}=\mathcal{B}_{1}\cup\mathcal{B}_{2}$ such that the runner-up is the IRV winner when using only the ballots in $\mathcal{B}_{1}$ and when using only the ballots in $\mathcal{B}_{2}$.
If an election is susceptible to both a plurality loser and runner-up paradox, we say the election is susceptible to a double reinforcement paradox. We now turn to the problem of determining if a ballot set $\mathcal{B}$ in a three-candidate election can be partitioned to produce a runner-up or plurality loser reinforcement paradox.
2. Necessary and Sufficient Conditions
To determine the likelihood that a three-candidate election is susceptible to a reinforcement paradox, we must be able to characterize when it is possible to partition $\mathcal{B}$ into two sub-elections $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ such that either the runner-up or the plurality loser wins both sub-elections.
Assume the preference profiles for complete and partial rankings are as shown in Tables 3 and 4 where in each case the expressions $X_{i},Y_{i}$, and $Z_{i}$ indicate the number of voters with the corresponding ranking. Note in Table 4 we do not list ballots of length two, as such ballots convey the same ranking information as the corresponding complete ballots.
Num. Voters $X_{1}$ $X_{2}$ $Y_{1}$ $Y_{2}$ $Z_{1}$ $Z_{2}$
1st choice AA AA $B$ $B$ $C$ $C$
2nd choice $B$ $C$ AA $C$ AA $B$
3rd choice $C$ $B$ $C$ AA $B$ AA
Table 3. Preference profile with complete preferences.
Num. Voters $X_{1}$ $X_{2}$ $X_{3}$ $Y_{1}$ $Y_{2}$ $Y_{3}$ $Z_{1}$ $Z_{2}$ $Z_{3}$
1st choice AA AA AA $B$ $B$ $B$ $C$ $C$ $C$
2nd choice $B$ $C$ AA $C$ AA $B$
3rd choice $C$ $B$ $C$ AA $B$ AA
Table 4. Preference profile with some partial preferences.
In the complete ballot case, let $X=X_{1}+X_{2}$, $Y=Y_{1}+Y_{2}$, and $Z=Z_{1}+Z_{2}$; in the partial ballot case, let $X^{\prime}=X_{1}+X_{2}+X_{3}$, $Y^{\prime}=Y_{1}+Y_{2}+Y_{3}$, and $Z^{\prime}=Z_{1}+Z_{2}+Z_{3}$. We assume without loss of generality that $A$ is the IRV winner and $C$ is the plurality loser.
First, we observe that if $A$ has a majority of first-place votes in the first round then the election is not susceptible to any kind of reinforcement paradox. That is, we cannot observe a paradox if $X>Y+Z$ (in the complete ballot case) or $X^{\prime}>Y^{\prime}+Z^{\prime}$ (in the partial ballot case). The reason is that if a $A$ has a majority, then they must earn a majority of first-place votes in either $\mathcal{B}_{1}$ or $\mathcal{B}_{2}$ for any partition of the ballot set.
Observation 1.
If an election contains a majority candidate then the election is not susceptible to a reinforcement paradox.
Thus, since $A$ is the IRV winner, we assume in the complete ballot case that
$Z<X,\ Z<Y\text{ and }Y+Z_{2}<X+Z_{1}$
and in the partial ballot case that
$Z^{\prime}<X^{\prime},\ Z^{\prime}<Y^{\prime}\text{ and }Y^{\prime}+Z_{2}<X^{\prime}+Z_{1}.$
In addition, as remarked previously, if there is a partition such that either the runner-up $B$ or the plurality loser $C$ wins both sub-elections in two rounds, then the order of elimination in $\mathcal{B}_{1}$ must be different from that in $\mathcal{B}_{2}$. (Suppose, for instance, that $B$ wins both sub-elections in two rounds and that $C$ is the plurality loser in both sub-elections. Using the notation of Table 1, this implies $X^{\prime}+Z_{1}<Y^{\prime}+Z_{2}$ in both sub-elections, which implies the same inequality holds in the original election.) This leads to the following.
Observation 2.
If there is a partition of the electorate in which either the runner-up $B$ or the plurality loser $C$ wins both sub-elections in the second round of counting, then the plurality loser is different in each sub-election.
As the following Propositions indicate, it is also possible for an election to have a partition in which the runner-up wins a majority of first-place votes in one sub-election while winning in round 2 in the other sub-election.
Proposition 1.
(Complete Ballots). An election demonstrates a runner-up reinforcement paradox if and only if the election satisfies
(1) $Y>X-Z_{2}+1\text{ and }Y>-X_{1}+X_{2}+Z_{1}-Z_{2}+1.$
That is, conditions (1) are necessary and sufficient for there to exist a partition of the election into two sub-elections in both of which the runner-up is the winner. If in addition,
(2) $Y>X+1\text{ and }Y>-X_{1}+X_{2}+Z+1,$
then there is a partition in which the runner-up wins a majority of first-place votes in one of the sub-elections and wins in the second round in the other sub-election.
An election demonstrates a plurality loser reinforcement paradox if and only if
(3) $Z$ $>X-Y_{2}+1,$
$Z$ $>X_{1}-X_{2}+Y_{1}-Y_{2}+1,$
$Z$ $>-X_{2}+Y+1,\text{ and }$
$Z$ $>\tfrac{1}{3}(X+Y)+1.$
In this case, there is a partition in which the plurality loser wins in the second round in both sub-elections.
Proposition 2.
(Partial Ballots). An election demonstrates a runner-up reinforcement paradox if and only if the election satisfies
(4) $Y^{\prime}$ $>X^{\prime}-Z_{2}+1,$
$Y^{\prime}$ $>-X_{1}+X_{2}+Z_{1}-Z_{2}+1,\text{ and}$
$Y^{\prime}$ $>\tfrac{1}{3}(-X_{1}+2X_{2}+2X_{3}+Z_{1}-Z_{2})+1.$
That is, conditions (4) are necessary and sufficient for there to exist a partition of the election into two sub-elections in both of which the runner-up is the winner. If in addition,
(5) $Y^{\prime}>X^{\prime}+1\text{ and }Y^{\prime}>-X_{1}+X_{2}+Z^{\prime}+1,$
then there is a partition in which the runner-up wins a majority of first-place votes in one of the sub-elections and wins in the second round in the other sub-election.
An election demonstrates a plurality loser reinforcement paradox if and only if
(6) $Z^{\prime}$ $>X^{\prime}-Y_{2}+1,$
$Z^{\prime}$ $>X_{1}-X_{2}+Y_{1}-Y_{2}+1,$
$Z^{\prime}$ $>-X_{2}+Y^{\prime}+1,\text{ and}$
$Z^{\prime}$ $>\tfrac{1}{3}(X^{\prime}+Y^{\prime})+1.$
In this case, there is a partition in which the plurality loser wins in the second round in both sub-elections.
In subsequent sections we sometimes consider a “limiting case” where the number of voters approaches infinity. In this case, the values at the top of a preference profile can be considered proportions rather than vote counts, in which case the “+1” term in the above propositions disappears. That is, the above propositions still hold for the limiting case by eliminating the plus 1 term, which corresponds to dividing 1 by a larger and larger number of voters as we take the limit.
The conditions in Proposition 2 reduce to those in Proposition 1 when $X_{3}=Y_{3}=Z_{3}=0$. The condition $Y^{\prime}>\frac{1}{3}(-X_{1}+2X_{2}+2X_{3}+Z_{1}-Z_{2})+1$ in Proposition 2 reduces to $Y>\frac{1}{3}(-X_{1}+2X_{2}+Z_{1}-Z_{2})+1$ and does not have an analogue in Proposition 1 because it is implied by the other conditions. To see this, note that
$X_{1}-2X_{2}$ $+3Y-Z_{1}+Z_{2}-3=(X-Y+Z_{1}-Z_{2})$
$+2(-X+Y+Z_{2}-1)+(X_{1}-X_{2}+Y-Z_{1}+Z_{2}-1)+X_{1}+(Y-Z).$
But $X+Y+Z_{2}-1\geq 0$ and $X_{1}-X_{2}+Y-Z_{1}+Z_{2}-1\geq 0$ by the remaining two conditions in (1), and YX>0Y-X>0 and $X-Y+Z_{1}-Z_{2}>0$ by assumption. Hence $X_{1}-2X_{2}+3Y-Z_{1}+Z_{2}-3>0$, which implies $Y>\frac{1}{3}(-X_{1}+2X_{2}+Z_{1}-Z_{2})+1$.
Before sketching the proof of the propositions, we make a few remarks. First, an election can be susceptible to either a runner-up or plurality loser reinforcement paradox when the IRV and plurality winners in the original election are distinct or when they coincide, for the complete ballot or partial ballot case.
Second, condition (2) clearly can only be satisfied if the IRV runner-up is the plurality winner. In this case, as the proof will demonstrate, if there is a partition in which the runner-up wins a majority of first-place votes in one sub-election, they must win in two rounds in the other sub-election with the original IRV winner being eliminated in round 1.
Third, it is easy to show that it is not possible to have a partition in which the plurality loser has a majority of first-place votes in one sub-election and wins in two rounds in the other sub-election. Instead, a plurality loser reinforcement paradox can only occur when the plurality loser wins in two rounds in both sub-elections.
We prove a portion of Proposition 1, as outlined below. The remaining proofs are analogous (and considerably longer).
Proof.
(of Proposition 1) We prove that the inequalities in (2) of Proposition 1 are necessary and sufficient for there to be a partition in which the runner-up wins a majority of first-place votes in one of the sub-elections and wins in the second round in the other sub-election.
Suppose the original election has a voter profile as shown in Table 3. Since candidate $B$ is not eliminated in either sub-election, we can collapse the two columns in Table 3 that rank $B$ first and simply consider the number of number of voters $Y$ who rank $B$ first and the remaining candidates in any order. Now suppose $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ are a partition of the election into two sets of ballots. Let the voter profile of $\mathcal{B}_{1}$ be as in Table 3 with $x_{1}$ replacing $X_{1}$, $x_{2}$ replacing $X_{2}$, $y$ replacing $Y$, $z_{1}$ replacing $Z_{1}$ and $z_{2}$ replacing $Z_{2}$, for some $0\leq x_{1}\leq X_{1}$, $0\leq x_{2}\leq X_{2}$, $0\leq y\leq Y,$ etc. Let $\mathcal{B}_{2}$ be the same except with $X_{1}-x_{1},$ $X_{2}-x_{2}$, YyY-y, $Z_{1}-z_{1}$ and $Z_{2}-z_{2}$.
Suppose $B$ wins by a majority of first-place votes in $\mathcal{B}_{2}$. Clearly, it also cannot win by a majority in $\mathcal{B}_{1}$ (otherwise $B$ would have a majority of first-place votes in the original election). So $B$ must win in two rounds in $\mathcal{B}_{1}.$ We claim that $A$ must be the plurality loser in $\mathcal{B}_{1}.$
To see this, suppose $C$ is the plurality loser in $\mathcal{B}_{1}.$ After $C$ is eliminated,
$x_{1}+x_{1}+z_{1}<y+z_{2}.$
On the other hand, since $B$ wins in a majority in $\mathcal{B}_{2},$
$(X_{1}-x_{1})+(X_{2}-x_{2})+(Z_{1}-z_{1})+(Z_{2}-z_{2})<Y-y.$
Solving both of these inequalities for $z_{1}$, and recalling that $z_{1}$ is an integer, we have
$X_{1}+X_{2}-Y+Z_{1}+Z_{2}-x_{1}-x_{2}+y-z_{2}+1\leq z_{1}\leq-x_{1}-x_{2}+y+z_{2}-1,$
which implies
$X_{1}+X_{2}-Y+Z_{1}+Z_{2}+2\leq 2z_{2}.$
Since $z_{2}\leq Z_{2}$, this implies $X_{1}+X_{2}-Y+Z_{1}+Z_{2}+2\leq 2Z_{2}$ or
$X_{1}+X_{2}+Z_{1}+2\leq Y+Z_{2},$
which is a contradiction, since $A$ is the IRV winner in the original election. This proves the claim.
Thus $A$ must be the plurality loser in $\mathcal{B}_{1}$ with $B$ winning in round 2. Let $z=z_{1}+z_{2}$ be the number of voters in $\mathcal{B}_{1}$ who first-rank $C$ and let $Z-z=(Z_{1}-z_{1})+(Z_{2}-z_{2})$ be the number of analogous voters in $\mathcal{B}_{2}$. We require (recalling that $X=X_{1}+X_{2}$)
(7) $x_{1}+x_{2}$ $<y$
(8) $x_{1}+x_{2}$ $<z$
(9) $x_{2}+z$ $<y+x_{1}$
(10) $(X-x_{1}-x_{2})+(Z-z)$ $<(Y-y).$
We simplify these inequalities in a series of steps in which we sequentially eliminate variables. In the first step, we collect lower and upper bounds for $y$.
Step 1: Solving (7), (8) and (10) for $y$ and adding the boundary constraints $0\leq y\leq Y,$ we have
$\max\{x_{1}+x_{2}+1,-x_{1}+x_{2}+z+1\}\leq y\leq\min\{Y,-X+Y-Z+x_{1}+x_{2}+z-1\}$
Since each lower bound must be bounded above by each upper bound, this implies
$x_{1}+x_{2}+1$ $\leq Y$ $\Rightarrow$ $x_{2}\leq Y-x_{1}-1$
$x_{1}+x_{2}+1$ $\leq-X+Y-Z+x_{1}+x_{2}+z-1$ $\Rightarrow$ $z\geq X-Y+Z+2$
$-x_{1}+x_{2}+z+1$ $\leq Y$ $\Rightarrow$ $z\leq Y+x_{1}-x_{2}-1$
$-x_{1}+x_{2}+z+1$ $\leq-X+Y-Z+x_{1}+x_{2}+z-1$ $\Rightarrow$ $x_{1}\geq\frac{1}{2}(X-Y+Z+2).$
Step 2: Collecting all the bounds (including (8)) on $z$, along with the boundary constraint $0\leq z\leq Z$, we have
$\max\{x_{1}+x_{2}+1,X-Y+Z+2\}\leq z\leq\min\{Z,Y+x_{1}-x_{2}-1\}.$
Again, since each lower bound must be bounded above by each upper bound, we have
$x_{1}+x_{2}+1$ $\leq Z$ $\Rightarrow$ $x_{2}\leq Z-x_{1}-1$
$x_{1}+x_{2}+1$ $\leq Y+x_{1}-x_{2}-1$ $\Rightarrow$ $x_{2}\leq\frac{1}{2}(Y-2)$
$X-Y+Z+2$ $\leq Z$ $\Rightarrow$ $Y\geq X+2$
$X-Y+Z+2$ $\leq Y+x_{1}-x_{2}-1$ $\Rightarrow$ $x_{2}\leq-X+2Y-Z+x_{1}-3$
Step 3: Collecting all the bounds on $x_{2}$, we have
$\max\{0\}\leq x_{2}\leq\min\{X_{2},Y-x_{1}-1,Z-x_{1}-1,\frac{1}{2}(Y-2),-X+2Y-Z+x_{1}-3\}.$
Hence
$0$ $\leq Y-x_{1}-1$ $\Rightarrow$ $x_{1}\leq Y-1$
$0$ $\leq Z-x_{1}-1$ $\Rightarrow$ $x_{1}\leq Z-1$
$0$ $\leq\frac{1}{2}(Y-2)$ $\Rightarrow$ $Y\geq 2$
$0$ $\leq-X+2Y-Z+x_{1}-3$ $\Rightarrow$ $x_{1}\geq X-2Y+Z+3$
Step 4: Finally, collecting all the bounds on $x_{1}$, we have
$\max\{0,\frac{1}{2}(X-Y+Z+2),X-2Y+Z+3\}\leq x_{1}\leq\min\{X_{1},Z-1,\}.$
This implies
$0$ $\leq Z-1$ $\Rightarrow$ $Z\geq 1$
$\frac{1}{2}(X-Y+Z+2)$ $\leq X_{1}$ $\Rightarrow$ $Y\geq-X_{1}+X_{2}+Z+2$
$\frac{1}{2}(X-Y+Z+2)$ $\leq Z-1$ $\Rightarrow$ $Y\geq X-Z+4$
$X-2Y+Z+3$ $\leq X_{1}$ $\Rightarrow$ $Y\geq\frac{1}{2}(X+Z+3)$
$X-2Y+Z+3$ $\leq Z-1$ $\Rightarrow$ $Y\geq\frac{1}{2}(X+4)$
Step 5: Collecting all the inequalities relating $X_{1}$, $X_{2}$, $Y$ and $Z$, these reduce to (assuming $Z\geq 4$):
$Y\geq X+2\quad\text{and}\quad Y\geq-X_{1}+X_{2}+Z+2.$
Since $Y$ is an integer, these conditions are equivalent to the conditions $Y>X+1$ and $Y>-X_{1}+X_{2}+Z+1$ in (2) of Proposition 1. ∎
Propositions 1 and 2 identify the necessary conditions on the original election for a runner-up or plurality loser reinforcement paradox to occur but do not identify any partitions. However, the inequalities that arise at each step of the propositions’ proofs can be used to identify election partitions leading to the paradoxes.
For instance, from the proof above of the necessity of (2) of Proposition 1, we have
  1. (1)
    $\max\{x_{1}+x_{2}+1,-x_{1}+x_{2}+z+1\}\leq y\leq\min\{Y,-X+Y-Z+x_{1}+x_{2}+z-1\}$
  2. (2)
    $\max\{x_{1}+x_{2}+1,X-Y+Z+2\}\leq z\leq\min\{Z,Y+x_{1}-x_{2}-1\}$
  3. (3)
    $\max\{0\}\leq x_{2}\leq\min\{X_{2},Y-x_{1}-1,Z-x_{1}-1,\frac{1}{2}(Y-2),-X+2Y-Z+x_{1}-3\}$
  4. (4)
    $\max\{0,\frac{1}{2}(X-Y+Z+2),X-2Y+Z+3\}\leq x_{1}\leq\min\{X_{1},Z-1,\}.$
Thus, if $X_{1}=50,$ $X_{2}=50$, $Y=102$, $Z_{1}=50$ and $Z_{2}=45,$ we can select (among other possibilities), a partition in which the first sub-election is
$x_{1}=48,x_{2}=40,y=89\mbox{ and }z=95,$
and the second partition is
$X_{1}-x_{1}=2,X_{2}-x_{2}=10,Y-y=13,Z-z=0.$
It is easily seen that $A$ wins the original election, but in $\mathcal{B}_{1}$ candidate $B$ wins in two rounds after $A$ is eliminated and in $\mathcal{B}_{2}$, candidate $B$ has a majority.
2.1. Consequences of the Algebraic Conditions
The inequalities identified in Propositions 1 and 2 allow us to identify further anomalies. For instance, Table 5 demonstrates that it is possible for a candidate to be the IRV winner, a Condorcet winner and a plurality winner and yet the profile still satisfies the requirements in Propositions 1 for both a runner-up and plurality loser paradox.
Num. Voters 1001 50 521 480 561 439
1st AA AA $B$ $B$ $C$ $C$
2nd $B$ $C$ AA $C$ AA $B$
3rd $C$ $B$ $C$ AA $B$ AA
Table 5. (Top) A profile with the IRV, plurality, and Condorcet winners coinciding that demonstrates both runner-up and plurality loser paradoxes.
The next example shows that reinforcement paradoxes can be nested: an election may exhibit a runner-up or plurality loser paradox and each sub-election demonstrate an additional reinforcement paradoxes.
Example 3.
Consider the profile in Table 6 where $A$ is the IRV winner. The election demonstrates a runner-up paradox and $B$ wins in both sub-elections $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$. However each of these sub-elections themselves satisfies the conditions for a runner-up paradox. This means that sub-election $\mathcal{B}_{1}$, in which $C$ is the plurality loser and $A$ is the runner-up, can be further subdivided into two smaller sub-elections in which $A$ is the IRV winner. (One such partition is $\mathcal{B}_{11}$ which corresponds to $X_{1}=12$, $X_{2}=31$, $Y_{1}=0$, $Y_{2}=25$, $Z_{1}=0$ and $Z_{2}=17$ and $\mathcal{B}_{12}$ which corresponds to $X_{1}=28$, $X_{2}=31$, $Y_{1}=5$$Y_{2}=26$, $Z_{1}=1$ and $Z_{2}=31$.) And sub-election $\mathcal{B}_{2}$, in which $A$ is the plurality loser and $C$ is the runner-up, can be further subdivided into two smaller sub-elections in which $C$ is the IRV winner. (One such partition is $\mathcal{B}_{21}$ which corresponds to $X_{1}=50$, $X_{2}=0$, $Y_{1}=4$, $Y_{2}=25$, $Z_{1}=30$ and $Z_{2}=0$ and $\mathcal{B}_{22}$ which corresponds to $X_{1}=10$, $X_{2}=38$, $Y_{1}=41$, $Y_{2}=54$, $Z_{1}=69$ and $Z_{2}=2$.)
Num. Voters 100 100 50 130 100 50
1st choice AA AA $B$ $B$ $C$ $C$
2nd choice $B$ $C$ AA $C$ AA $B$
3rd choice $C$ $B$ $C$ AA $B$ AA
Partition $\mathcal{B}_{1}$ 40 62 5 51 1 48
$\mathcal{B}_{2}$ 60 38 45 79 99 2
Table 6. A profile that demonstrates a runner-up paradox with sub-elections that demonstrate additional reinforcement paradoxes.
Additional instances of the paradox may occur in which an election can be divided into three or more sub-elections in which either the runner-up or plurality loser wins.
Num. Voters 1000 999 560 438
1st choice AA $B$ $C$ $C$
2nd choice $C$ $C$ AA $B$
3rd choice $B$ AA $B$ AA
Table 7. A profile that demonstrates a plurality loser, but not a runner-up, reinforcement paradox.
Finally, we note that elections may be susceptible to a plurality loser paradox but not a runner-up paradox, as demonstrated by the profile in Table 7 where neither Condition (1) nor (2) of Proposition 1 is satisfied, but the plurality loser inequalities are satisfied. Such preference profiles can arise when voter preferences are “single-peaked.” In this case, candidate $C$ is like a “centrist” candidate in a spatial model, as if $A$, $B$, and $C$ have been arranged on a left-to-right line, with $C$ in the middle.
Of course, real-world elections typically do not return ballot data resembling the profile in Table 7. In fact as we will see in Section 5 in real elections, the IRV winner tends to be a much stronger candidate than the plurality loser. Thus, we expect that if an election is susceptible to a plurality loser reinforcement paradox in this data then it will also be susceptible to a runner-up paradox.
3. Threshold Results
In this section we examine the minimal percentage each candidate can earn while the election still demonstrates a reinforcement paradox. For some classical paradoxes such as upward and downward monotonicity paradoxes, an election can exhibit such a paradox only if each of the three candidates is not “weak.” We show this is not the case for a general reinforcement paradox in Proposition 3, where the plurality loser can have an arbitrarily low percentage of the first-places votes and yet an election is susceptible to a (runner-up) reinforcement paradox. We show in Proposition 4 that for an election to be susceptible to a runner-up (respectively plurality loser) paradox, the IRV runner-up (respectively plurality loser) must earn at least 25% of the first-place votes.
In this section, we assume that the number of voters approaches infinity, so that we use proportions of votes rather than integers. We also assume that each voter provides a complete ranking as this does not affect the logic of the proofs. Since the number of voters approaches infinity, we ignore the “+1” which appears in the Propositions in Section 2, as this term limits to zero.
Proposition 3.
For any $\epsilon>0$, there is an election which is susceptible to the reinforcement paradox and the first-place vote share for the plurality loser is less than $\epsilon$.
Proof.
Consider the profile in Table 8, where $p_{A}$ is the proportion of voters who cast the ballot $A\succ B\succ C$ and also the proportion of voters who cast the ballot $A\succ C\succ B$, and $\beta$ and $\delta$ are positive and very small. Furthermore, assume $2p_{A}+\beta<0.5$ and $4p_{A}+\beta+\delta=1$ (which implies $\delta>\beta)$. Note that $A$ is the IRV winner and $C$ is the plurality loser.
This profile satisfies condition (2) of Proposition 1 and thus the election demonstrates a runner-up reinforcement paradox. Since $\delta$ is arbitrary, we can make $\delta<\epsilon$ and the proposition is proven. For completeness, we include a partition of the electorate into two sub-electorates, as shown in the table. Let $\delta_{B}$ and $\delta_{A}$ be positive small parameters such that $0<\delta_{B}<\delta_{A}<\delta/2$. Then $B$ wins when using the ballots of both sub-electorates separately. ∎
Prop. Voters $p_{A}$ $p_{A}$ $2p_{A}+\beta$ $\delta$
1st choice AA AA $B$ $C$
2nd choice $B$ $C$ AA AA
3rd choice $C$ $B$ $C$ $B$
Partition $\mathcal{B}_{1}$ $\delta-\delta_{A}$ 0 $\delta-\delta_{B}$ $\delta$
$\mathcal{B}_{2}$ $p_{A}-\delta+\delta_{A}$ $p_{A}$ $2p_{A}+\beta-\delta+\delta_{B}$ 0
Table 8. A profile in which the plurality loser earns a small proportion of first-place votes and the election demonstrates a (runner-up) reinforcement paradox.
Thus, for a large enough electorate the plurality loser can control an arbitrarily small amount of first-place votes and the election can demonstrate a reinforcement paradox. This proposition distinguishes reinforcement paradoxes from other classical paradoxes such as upward or downward monotonicity paradoxes. For a three-candidate election to demonstrate an upward (resp. downward) paradox, each candidate must earn at least 25% (respectively $16\frac{2}{3}\%$) of the first-place votes [39]. That is, for some classical paradoxes none of the candidates can have an extremely weak first-place showing; each candidate must be “viable” in some sense. Proposition 3 shows this is not the case for the reinforcement paradox. However, in the next proposition we show that for a runner-up (resp. plurality) reinforcement paradox, the runner-up (resp. plurality loser) must earn at least 25% of the first-place votes.
Proposition 4.
If an election is susceptible to a runner-up (resp. plurality loser) reinforcement paradox then the runner-up (resp. plurality loser) must earn at least 25% of the first-place votes in round 1. This bound is tight in the limit as the number of voters approaches $\infty$.
Proof.
Runner-up case. Recall that in order for an election to demonstrate a reinforcement paradox, no candidate can earn an initial majority, and thus candidate $A$ must earn less than 50% of the first-place votes. This implies $B$ must earn at least $25\%$ of the first-place vote.
To see that that bounds are tight, consider the profile in Table 9 where $\epsilon_{A},\epsilon_{B},\epsilon_{1}$ and $\epsilon_{2}\geq 0$ are small parameters satisfying $\epsilon_{B}=\epsilon_{A}-\epsilon_{1}+\epsilon_{2}.$ If $\epsilon_{B}>\epsilon_{2}-\epsilon_{A}$ then this profile satisfies the conditions in (2) of Proposition 1.
Num. Voters $0.50-\epsilon_{A}$ $0.25+\epsilon_{B}$ $\epsilon_{1}$ $0.25-\epsilon_{2}$
1st choice AA $B$ $C$ $C$
2nd choice $B$ AA AA $B$
3rd choice $C$ $C$ $B$ AA
Table 9. Minimum profile for runner-up reinforcement paradox
Plurality loser case. From Proposition 1, $Z>\frac{1}{3}(X+Y)$ and thus the plurality loser must have at least $25\%$ of the first-place vote. On the other hand, consider the profile in Table 10 in which the plurality loser has just over 25% of the vote and yet, for some suitably small $\epsilon>0,$ satisfies the plurality loser inequalities from Proposition 1. ∎
Num. Voters $0.375-\epsilon$ $0.375-2\epsilon$ $0.25+3\epsilon$
1st choice AA $B$ $C$
2nd choice $C$ $C$ AA
3rd choice $B$ AA $B$
Table 10. Minimum profile for plurality loser reinforcement paradox
We conclude this section by noting that there are elections susceptible to a reinforcement paradox in which one of the sub-elections has an arbitrarily small percentage of voters.
Example 4.
Let $m\in\mathbb{N}$ be a large even number. Consider an election $\mathcal{B}$ with 3m13m-1 ballots (complete or partial) in which the first and second preferences are distributed as follows: $m$ ballots of the form $A\succ B$; $m+1$ ballots of the form $B\succ A$; $m/2$ ballots of the form $C\succ A$; and m/22m/2-2 ballots of the form $C\succ B$. Candidate $A$ wins this election in two rounds after the $C$ is eliminated in round 1.
Let $\mathcal{B}_{1}$ consist of the sub-election with $9$ ballots: $2$ ballots of the form $A\succ B$; $3$ ballots of the form $B\succ A$; and $4$ ballots of the form $C\succ A$. Let $\mathcal{B}_{2}$ consist of the sub-election with 3m103m-10 ballots: m2m-2 ballots of the form $A\succ B$; m2m-2 ballots of the form $B\succ A$; m/24m/2-4 ballots of the form $C\succ A$; and m/22m/2-2 ballots of the form $C\succ B$. Then $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ form a partition of $\mathcal{B}$, and candidate $B$ wins both of these sub-elections. Since $m$ can be arbitrarily large, this implies there exists an election demonstrating a reinforcement paradox in which one of the sub-elections consists of an arbitrarily small fraction of the ballots in the original election.
Example 4 is interesting as it suggests that $B$ was “close” to being a winner in the original election. In fact, $B$ might feel it particularly unfair that the addition of a handful of ballots (from $\mathcal{B}_{1}$, where $B$ was a winner) caused $B$ to go from winner (in $\mathcal{B}_{2}$) to loser (in $\mathcal{B}_{1}\cup\mathcal{B}_{2}$).
4. Results under various models
In this section we use different models of voter behavior to generate millions of synthetic elections. Among the many possible models we could use, we choose models which have been used extensively in prior social choice research. For each model, we generate 100,000 profiles and check whether a generated election is susceptible to a reinforcement paradox. Thus, our primary tool for providing (estimated) probabilities is Monte Carlo simulation, although we can provide exact probabilities using other means for the impartial anonymous culture (IAC) model.
We are interested in probabilities for relatively large electorates, as the real-world elections we analyze in the next section have electorate sizes ranging from a few hundred to almost one million (although most of the elections contain 20,000 or fewer voters). Thus, our simulations use an electorate size of $V=1001$ voters. We also ran simulations with $V\in\{601,2003,3901\}$, but the estimated probabilities were not markedly different for these other choices of $V$. With the exception of the IAC model, in this section we consider probabilities only for the complete ballot case.
4.1. Descriptions of the Models
We use two different types of models: one-dimensional spatial models and models which sample over the ballot simplex. In the spatial framework, we consider three different models. The first model, which we denote 1D Spatial, places three candidates and $V=1001$ voters on the real line using the standard normal distribution, and a voter’s preference ranking is determined by their Euclidean distance from each candidate. For example, if the three candidates $A$, $B$, and $C$ are positioned by the normal distribution at 1.2-1.2, 0.1-0.1, and 1.1 respectively, then a voter positioned at 0.3 would cast a ballot with $B$ ranked first, $C$ ranked second, and $A$ ranked third. Such one-dimensional spatial models have been widely used in the social choice literature [17, 18, 29, 51].
Because of the increased political polarization in the United States and elsewhere, we also use two bimodal spatial models to investigate the potential effects of polarization on the probability of the reinforcement paradox. Under the second spatial model, which we denote 1D Bimodal, we create a bimodal distribution from an equally weighted combination of the normal distribution centered at 2-2 and the normal distribution centered at 2, both with standard deviation 1. See Figure 1 (top) for a visualization of this bimodal distribution. We then place candidates and voters along the line using the generated (normalized) bimodal distribution as a probability distribution, similar to 1D Spatial. Under the third spatial model, which we denote 1D Bimodal Weighted, we use the same two normal distributions to generate the bimodal distribution but give 50% more weight to the left distribution, resulting in a bimodal distribution as shown Figure 1(bottom). We use this model to capture polarization when one side of the political spectrum is stronger than the other. (We do not investigate the effect of changing the variance.)
Figure 1. Visualizations of the two bimodal spatial models.
[…]
If we fix an $\alpha$, we can consider the limiting probability of observing a paradox where we let $V\rightarrow\infty$. In this limit, the discrete ballot simplex becomes the continuous 5-simplex determined by $X_{1}+X_{2}+Y_{1}+Y_{2}+Z_{1}+Z_{2}=1$, where each term in the sum is represents a (non-negative) proportion of votes. By using large electorates in our simulations, we estimate limiting probabilities under each model.
4.2. Results
Throughout this section and the next, in our analysis we identify four probabilities: the probability that a runner-up (RU) paradox occurs; the probability that a plurality loser (PL) paradox occurs; the probability that at least one of these (RU or PL) paradoxes occurs; and the probability that both (RU and PL) paradoxes occur.
We begin by providing exact probabilities under Dir(1)(1), or IAC model. Because of its simplicity and prevalence in the social choice literature, researchers have dedicated significant effort to developing tools for calculating probabilities under this model. One of the fruits of these labors is the software package Normaliz [9], which can provide limiting probabilities under IAC in three-candidate elections given a set of linear constraints. In the complete ballot case, Normaliz determines the proportion of integer values of $X_{1},X_{2},Y_{1},Y_{2},Z_{1}$ and $Z_{2}$ that satisfy the inequalities in Proposition 1 given a fixed number of voters $V$, as $V\rightarrow\infty$. In the partial ballot case, Normaliz, does the same thing for $X_{1},X_{2},X_{3},Y_{1},Y_{2},Y_{3},Z_{1},Z_{2}$ and $Z_{3}$ that satisfy the inequalities in Proposition 2.
The probabilities in the following proposition are thus exact (rather than approximated through simulation) and are obtained directly from Normaliz by inputting the appropriate sets of inequalities. The conditional probabilities are obtained by adding an additional requirement to limit the number of elections to those without a majority candidate.
Proposition 5.
The limiting probabilities that an election is susceptible to a reinforcement paradox under the Dir(1)(1) model (the IAC model) are as shown in Table 11 for complete and partial ballots. The “Cond. Prob.” probabilities are conditioned on the absence of a majority candidate.
PL or RU RU PL PL and RU
Prob. $\frac{191}{768}\approx 24.9\%$ $\frac{127}{576}\approx 22.5\%$ $\frac{79}{1152}\approx 6.9\%$ $\frac{31}{768}\approx 4.0\%$
Cond. Prob. $\frac{191}{336}\approx 56.8\%$ $\frac{127}{252}\approx 47.2\%$ $\frac{79}{504}\approx 15.7\%$ $\frac{31}{336}\approx 9.2\%$
Prob. $\frac{1341257}{5308416}\approx 25.3\%$, $\frac{527}{2304}\approx 22.9\%$ $\frac{7031}{110592}\approx 6.4\%$ $\frac{210439}{5308416}\approx 4.0\%$
Cond. Prob. $\frac{1341257}{3006720}\approx 44.6\%$ $\frac{527}{1305}\approx 40.4\%$ $\frac{7031}{62640}\approx 11.2\%$ $\frac{210439}{3006720}\approx 7.0\%$
Table 11. Probabilities under the Dir(1)(1) (IAC) model. The top two rows correspond to the complete ballot case; the bottom two rows correspond to the partial ballot case.
The probabilities of one or more paradoxes occurring are similar in the complete and partial ballots cases. When conditioned on the absence of a majority candidate, the probabilities are significantly lower in the partial ballot case. This is because the probability an election does not contain a majority candidate is $145/256=56.64\%$ in the 8-simplex but is only $7/16=43.75\%$ in the 5-simplex. Overall, the probabilities are fairly high. This is largely because they indicate the percentage of elections in which there exists a partition into two sub-elections demonstrating a reinforcement paradox; the numbers do not indicate anything about the number of such partitions in any given election.
[…]
As $\alpha$ increases, all four probabilities also increase; this is unsurprising because as $\alpha$ increases, a typical generated election becomes “closer” or “more competitive.” When $\alpha$ is small, many generated elections represent landslide victories where one candidate has a strong majority of first-place votes. When $\alpha$ is large, the generated elections tend to cluster in the center of the simplex where there is no majority candidate. This also explains why we see a large increase in probabilities when moving from the top of Figure 2 to the bottom when $\alpha$ is small (when there are few elections without a majority candidate) but not when $\alpha$ is large (when almost all elections have no majority).
Table 12 provides specific results from the simulations for a few choices of $\alpha$. Recall that 100,000 simulations were run for each choice of $\alpha$, so that when $\alpha=0.1$, for instance, only 3.5% of generated elections contained no majority candidate. Note that we include simulations for $\alpha=1$ even though we have exact probabilities from Proposition 5.
Table 12 also displays information for three large $\alpha$ values which are not shown in Figure 2. Recall that as $\alpha\rightarrow\infty$ the distribution resembles the IC model, in which each voter independently chooses one of the six possible rankings at random. For comparison, we also conducted simulations using the IC model with $V=1001$, which returned an estimated probability of 1.0 for the runner-up, plurality, or both (RU, PL, and RU and PL) paradoxes occurring. Intuitively, this is clear since when every voter chooses a ranking at random, $X_{1}\approx X_{2}\approx Y_{1}\approx Y_{2}\approx Z_{1}\approx Z_{2}$. Thus, in the limit, Condition (2) and the plurality loser inequalities of Proposition 1 will always be satisfied. It is unsurprising, therefore, that the probabilities in Figure 2 appear to be trending toward 1 and the values at the bottom of Table 12 indicate convergence to 1 in all cases.
$\alpha$ PL or RU RU PL PL and RU No Maj. Cand.
0.1 1257 1033 242 18 3500
0.5 12419 10737 2730 1048 24024
1.0 24745 21852 6825 3932 43590
1.5 34319 31065 11082 7828 56455
2.0 42038 38548 15426 11936 65795
2.5 48079 44538 19721 16180 72663
3.0 53702 50123 23732 20153 78324
10.0 85618 83538 62924 60844 98498
20.0 96383 95392 86120 85219 99948
30.0 98814 98442 94367 93955 99997
$\infty$ (IC) 100000 100000 100000 100000 100000
Table 12. Simulation results for some choices of $\alpha$. For each $\alpha$ we generated 100,000 profiles.
The 1D spatial models provide another way to estimate the probabilities of a reinforcement paradox, in part because the preferences they represent are single-peaked. Thus, in simulations in which the positions of the candidates $A$, $B$ and $C$ are ordered left to right, for instance, $X_{2}=Z_{1}=0$ since $B$ must be either first or second-ranked by all voters.
The results under all spatial models are shown in Table 13. The first row gives unconditioned probabilities; the second row gives probabilities conditioned on the absence of a majority candidate for 1D spatial. The middle (respectively last) two rows give the corresponding probabilities under 1D Bimodal (respectively 1D Bimodal Weighted).
Interestingly, the corresponding probabilities under 1D Spatial and 1D Bimodal are approximately equal, perhaps because both models are symmetric about zero. When we unbalance the distribution under 1D Bimodal Weighted, the probabilities change significantly. In particular, 1D Bimodal Weighted is the only model in this study under which it is more likely to observe a plurality loser paradox than a runner-up paradox. We do not know with certainty the reason for this, but we speculate that it is due to different responses to the so-called “center squeeze” [49]. In one-dimensional three-candidate spatial models, the central candidate will often be the plurality loser even if they are closest to the median voter. Using the notation of Section 2, if $C$ is the central candidate then $X_{1}=Y_{1}=0$ since $C$ must be second-ranked by all voters. Hence the conditions in Case (1) of Proposition (1) cannot be satisfied (since they imply $B$ has the majority of the first-place vote). The conditions in Case (2) reduce to
$Y>X-Z_{2}\quad\mbox{ and }\quad Y>X_{2}+Z_{1}-Z_{2}=X+Z_{1}-Z_{2}.$
If the distribution of voters is symmetric, $Z_{1}$ and $Z_{2}$ are likely to be similarly-sized, and so the rightmost inequality will be satisfied whenever $Y>X$. Thus, a runner-up (RU) paradox will occur often when the IRV and plurality winners are distinct.
If the distribution of voters is not symmetric, as with the 1D Bimodal Weighted model, the left and right-wing candidates are likely to be near the left and right peaks of the distribution respectively, with the centrist in between but on the “upward slope” of the larger peak (the left one in Figure 1). Thus, the quantity $Z_{1}-Z_{2}$ is more likely to be larger than in the symmetric case, decreasing the likelihood of Case (2) occurring. A similar argument holds if $A$ or $B$ is the central candidate.
For the plurality loser (PL) paradox to occur, it is easily seen that $C$ must be the central candidate. In this case, $C$ is again likely to be on the upward slope of the larger peak in the 1D Bimodal Weighted model (and hence $Z$ is likely to be larger) than in the 1D Bimodal model. So the probability of the plurality loser (PL) paradox occurring is greater in the nonsymmetric case than the symmetric.
Under all three 1D spatial models, the probability of a double paradox (both RU and PL) is extremely small as it is rare for both $Y$ and $Z$ to be sufficiently large simultaneously.
PL or RU RU PL PL and RU
Unconditioned Probability 16.63 12.17 4.47 0.01
Given No Majority Cand. 40.07 29.32 10.76 0.01
Unconditioned Probability 16.68 12.52 4.18 0.02
Given No Majority Cand. 34.05 25.55 8.53 0.03
Unconditioned Probability 11.46 5.05 6.46 0.05
Given No Majority Cand. 26.48 11.67 14.92 0.12
Table 13. Estimated probabilities, reported as percentages, under our three spatial models. The top two rows give probabilities for 1D Spatial, the middle two rows for 1D Bimodal, and the last two for 1D Bimodal Weighted.
5. Empirical Results
In this section we provide empirical results using a large database of real-world IRV elections. Our data comes from three sources. First, we use single-winner political elections from the United States with at least three (not write-in) candidates. This data is available at the FairVote data repository [44], and contains ballot data from municipal elections in cities such as San Francisco, CA and Minneapolis, MN. The repository also has ballot data for federal elections in the states of Alaska and Maine. Second, we use single-winner elections from the American Psychological Association (APA), which generally is willing to share ballot data for ranked-choice APA presidential elections and elections for the Board of Directors. Some of this data is available on preflib.org [33], and some was shared directly with the first author. Third, we use single-winner political elections from a database of Scottish local government elections. The ballot data for these elections is available at https://github.com/mggg/scot-elex.
In total, we accessed the preference profiles for 352 single-winner IRV elections with at least three candidates. [^3] Any reference to the number of candidates in an election does not include write-ins. In our dataset, write-in candidates receive an insignificant amount of votes and ignoring them does not affect the analysis. Many of these elections contain more than three candidates; because we are interested in the three-candidate case, for each election we “reduce” the number of candidates by running the IRV algorithm until only three candidates remain and then analyze the resulting profile. Note that the relationship between the original election and the resulting three-candidate election is not clear with respect to the susceptibility to reinforcement paradoxes. It is a priori possible that the original election demonstrates a paradox while the reduced three-candidate election does not, and vice versa. We use this methodology because we are interested in three-candidate elections and analyzing the elections with more than three candidates using the IRV algorithm in this way allows us to expand the size of our three-candidate election dataset. Moreover, in some cases if the resulting three-candidate is susceptible to a paradox then we can gain insight into how the original election might be susceptible to a paradox; this is how a reinforcement paradox in a five-candidate Minneapolis city council election was found, for example [35].
Jurisdiction USA APA Scotland Total
Elections 295 27 30 352
3-Cand. Elections 104 10 1 115
3-Cand. Elections, no Majority Cand. 44 6 1 51
3-Cand. Reduced Elections, no Maj. Cand. 164 20 24 208
Table 14. A summary of our data sources.
5.1. Results
Table 15 provides the results for elections having three candidates in the original election and for all 352 elections. We use the same abbreviations RU and PL as in Section 4. The first row indicates the number and percentage of elections exhibiting one or more paradoxes among the three-candidate elections. For example, of the 115 elections which originally contained three candidates, 17 of them (14.8%) exhibit a reinforcement paradox. The second row indicates the same information but for all the elections including the reduced elections. The third and fourth rows indicate the percentage of elections that exhibit the given type of paradox, conditioned on the absence of a majority candidate.
Unlike the simulated elections in Section 4, all the elections in the database that demonstrate a plurality loser paradox also demonstrates a runner-up paradox. (Note that the two rightmost columns of Table 15 are identical.) This is because in real elections runner-up candidates tend to be much stronger than plurality loser candidates.
PL or RU RU PL PL and RU
3-Cand. Elections 17, 14.8% 17, 14.8% 7, 6.1% 7, 6.1%
3-Cand. Reduced Elections 86, 24.4% 86, 24.4% 27, 7.7% 27, 7.7%
3-Cand. Elections, no Maj. Cand. 33.3% 33.3% 13.7% 13.7%
3-Cand Reduced Elections, no Maj. Cand. 41.3% 41.3% 13.0% 13.0%
Table 15. Reinforcement paradox results for the 115 real-world three-candidate elections and for all 352 “reduced” elections.
In Section 3, we showed in Proposition 3 that if an election demonstrates a runner-up reinforcement paradox then the runner-up could receive as little as 25% of the first-place votes and the plurality loser could receive an arbitrarily small percentage of the first-place votes. Similarly, if an election demonstrates a plurality loser paradox then the plurality loser can receive as little as 25% of the first-place votes. How close to these minima do we come in real-world elections? The lowest percentage of first-place support earned by the plurality loser in an election which demonstrated a reinforcement paradox is 3.6%, which occurred in the 2023 city council ward 7 election in Minneapolis. Table 16 shows the preference profile for this election, which originally contained three (not write-in) candidates; that is, the profile displayed is not obtained from eliminating previous official candidates. The three candidates were Cashman ($C$), Graham ($G$), and Foxworth ($F$). Foxworth earns only 280 (3.6%) first-place votes of the 7969 ballots cast, yet the runner-up inequalities in Condition (2) of Proposition 2 are satisfied.
Num. Voters 1440 1217 1212 1473 611 1726 185 68 37
1st $C$ $C$ $C$ $G$ $G$ $G$ $F$ $F$ $F$
2nd $G$ $F$ $C$ $F$ $C$ $G$
3rd $F$ $G$ $F$ $G$ $G$ $C$
Table 16. The preference profile for the 2023 city council ward 7 election in Minneapolis.
This 2023 Minneapolis city council election is an interesting example of a reinforcement paradox, even though it initially appears unremarkable from an IRV perspective. In this election the plurality, Condorcet, and IRV winners are identical, ruling out the possibility of classic issues like a monotonicity paradox, spoiler effect, or similar pathology. While the election is close—Cashman and Graham receive a comparable number of first-place votes—Cashman is the clear winner. From a social choice perspective, this election typically warrants no further analysis. However, the presence of a reinforcement paradox makes this election more compelling than it might appear at first glance.
The lowest first-place vote percentage earned by the IRV runner-up in an election which demonstrates a runner-up paradox is 28.9%, achieved in the 2018 mayoral election in San Francisco. This election had 9 official candidates, and thus the 28.9% represents the vote share of the runner-up after eliminating six candidates. The lowest percentage for a plurality loser in an election which demonstrates a plurality loser election is 26.9%, achieved in the 2008 County Executive election in Pierce County, Washington. This election had four official candidates, and the plurality loser earned 26.9% of the voter after eliminating one candidate.
5.2. Bootstrapping Results
In Section 4 we ran simulations based on sampling using theoretical models such as the IAC model. Since we have access to a large dataset of real-world elections, we can also sample from these elections and run simulations using nonparametric bootstrap analysis.
Bootstrapping allows us to generate random samples from each election profile to create a set of “pseudoprofiles.” Analysis of these pseudoprofiles can give a sense of how “close” the given election is to demonstrating a paradox. For instance, if we sample from an election profile that does not demonstrate a reinforcement paradox and determine that none of the sampled pseudoprofiles demonstrates a paradox, then the original election is not close to an election that demonstrates a paradox. On the other hand, if an election demonstrates a paradox but relatively few of its pseudoprofiles do, then the election is close to one which does not demonstrate a paradox. This is the case, for example, in the 2023 city council ward 7 election in Minneapolis (Table 16). The election demonstrates a reinforcement paradox but only 3370 of 10000 generated pseudoprofiles demonstrate a paradox. Thus, we could argue that this election demonstrating a paradox is not the most likely outcome.
To perform this analysis, we follow techniques used in the classical papers [48] and [50]. For each of the 352 three-candidate profiles from our real-world dataset, we created 10,000 pseudoprofiles by sampling ballots with replacement from the actual three-candidate profile. This sampling allowed us to generate 3,520,000 three-candidate elections which we can check for reinforcement paradoxes.
To keep the computation time reasonable and to match the analysis in Section 4, if the actual profile contained more than 1,001 voters then for each generated pseudoprofile we sampled only 1,001 ballots; otherwise each pseudoprofile contained the same number of voters as the original election. Most of the elections in the dataset contain more than 1,001 voters and thus most of the generated elections contain 1,001 voters.
Table 17 shows the probability of each type of reinforcement paradox across the 3,520,000 pseudoprofiles. The results are similar to the “3-Cand Reduced Elections” probabilities in Table 15, and thus the bootstrapping results do not meaningfully affect how frequently we expect such paradoxes to occur in real-world elections.
A more nuanced summary of the bootstrapping results can be seen in Figure 3. Each histogram illustrates the frequency of the proportion of pseudoprofiles generated by each original election that resulted in a “paradoxical” election. The top image is based on the 86 original three-candidate elections that were susceptible to the paradox. Thus, for example, the right-most bar on the top histogram indicates that in 33 of these elections, between 95 and 100% of the associated pseudoprofiles were susceptible to the paradox. Similarly, the second-to-the-right bar indicates that in 10 of these elections, between 90 and 95% of the associated pseudoprofiles were susceptible to the paradox. The bottom histogram is constructed the same way for the 266 original three-candidate elections which were not susceptible to the paradox with the exception that elections that generated zero “paradoxical” pseudoprofiles are omitted for clarity (since the left bar would be out-sized).
From Figure 3 (top), we can observe that roughly half of the original elections that were susceptible to a paradox were extremely likely to exhibit that paradox in the sense that more than 90% of their associated pseudoprofiles also did so. Most of the remaining elections in this category also generated pseudoprofiles that more often than not were susceptible to a reinforcement paradox. The reverse is true in Figure 3 (bottom): the large majority of the original elections that were not susceptible to the paradox generated few pseudoprofiles that were susceptible to a reinforcement paradox.
Noteworthy, however, is the fact that across all 352 original elections, 206 generated at least one pseudoprofile which was susceptible to a reinforcement paradox. Thus, arguably, 206 elections are close to an election that is susceptible to a paradox. If we restrict the analysis to elections that did not contain a majority candidate, 187 of the 208 of the original elections generated at least one pseudoprofile that is susceptible to a reinforcement paradox. These results suggest that almost any three-candidate real-world election which does not contain a majority candidate is “close” to an election that is susceptible to a reinforcement paradox. We note such a high percentage is not achieved under bootstrapping for other paradoxes. For example, in [37] the same type of analysis was performed for the spoiler effect using a smaller dataset of American IRV elections (but still using 10,000 pseudoprofiles per election), and only five elections generated at least one pseudoprofile demonstrating a spoiler effect under IRV. Thus, it is not the case that performing bootstrap analysis guarantees that almost all elections in a dataset will be “close” to an election that exhibits a given paradox.
Finally, recall in the 352 actual elections, any election that was susceptible to a plurality loser paradox was also susceptible to a runner-up paradox. This is not the case for pseudoprofiles: 9,542 generated elections are susceptible to a plurality loser but not to a runner-up paradox. Thus, while such an outcome is not likely, some real-world elections are close to this phenomenon occurring.
PL or RU RU PL PL and RU
Unconditioned Prob. 24.21% 23.94% 7.21% 6.94%
Conditioned Prob. 42.29% 41.82% 12.60% 12.13%
Table 17. Probability, expressed as a percentage, of observing a reinforcement paradox in the 3,520,000 generated pseudoprofiles.
Figure 3. (Top) Bootstrapping frequency results for elections that are susceptible to a reinforcement paradox. (Bottom) Bootstrapping frequency results for elections that are are not susceptible to a reinforcement paradox but generated at least one pseudoprofile that is.
6. Discussion
In this section we discuss our results, starting with a comparison between the theoretical and empirical results, then placing our results into context by comparing them to previous results on the reinforcement paradox.
6.1. Theoretical Versus Empirical Results
As is often the case in social choice, none of the theoretical models give probabilities which closely match what we observe in real-world elections across all four paradox types. Interestingly, if we focus on the 352 three-candidate reduced elections (lines two and four of Table 15), the IAC complete-ballot model (Table 11) provides a decent approximation of the empirical results. Furthermore, even though the real-world elections contain many partial ballots, generally the complete-ballot IAC model more accurately reflects the empirical probabilities across the 352 reduced elections than the partial-ballot IAC model. For other voting paradoxes such as monotonicity paradoxes or the spoiler effect, the IAC model typically gives probabilities much higher than what we observe in real elections. If we focus only on the 151 three-candidate actual elections, Dir(0.6)(0.6) gives the best approximation of the empirical unconditioned results (with the exception of “PL and RU”), which makes sense because in these elections, the winner is stronger overall than in the Dir(1)(1) (IAC) model, consistent with the election data. Overall, the 1D-spatial models, both normal and bimodal, underestimate the probabilities of one or more types of paradox occurring.
What none of the models capture is that in actual elections, it is extremely rare to observe a plurality loser paradox without also observing a runner-up paradox. Under spatial models this difference is especially pronounced, as under those models if a plurality loser paradox is exhibited then we almost never also observe a runner-up paradox. This is generally true across all models, even though the runner-up (RU) paradox is roughly three times as likely to occur as the plurality loser (PL) paradox.
All theoretical models predict that if an election is somewhat “competitive” then the probability of demonstrating a reinforcement paradox is large (although usually not larger than 50%). The empirical results also bear this out, with empirical probabilities sometimes being larger than theoretical ones, even when using theoretical models which often produce close elections. Of course, the empirical probabilities come nowhere close to the probability of 1 predicted by IC, the model which tends to produce extremely competitive elections.
6.2. Comparison to Previous Reinforcement Paradox Results
As mentioned in the Introduction, we are aware of only two previous papers [12, 47] which estimate the probability with which an election demonstrates a reinforcement paradox. Our work differs from theirs in several significant ways. Most importantly, in most cases the analysis in these earlier papers was based on simulating two separate elections, checking whether the elections had the same winner, and then indicating a reinforcement paradox had occurred if the winner obtained from the combined election was different from the winner in the two generated elections. That is, the authors did not start with one election and look for a partition of the ballots as we do here. As Plassmann and Tideman state in [47], “We determine the frequencies of the reinforcement paradox by sampling ballots for two separate elections, which we then combine to determine whether the paradox occurred for any voting rule. In contrast, we determine the frequencies of all other paradoxes from a single sample of ballots.” Plassmann and Tideman do not try to decompose one set of ballots into a partition as we do, stating “There is no obvious way in which to divide the ballots from a single election into two distinct sets that would yield the appropriate distribution of rankings in the two new sets and would permit us to examine reinforcement.” Courtin et al. [12] generally take a similar approach, simulating two ballot sets and then combining them. We have contributed to the investigation of this paradox by identifying a method in the 3-candidate case, through Propositions 1 and 2, to determine if such a division is possible, without having to separately test each possible partition of a ballot set into two sub-elections. We feel our approach is equally valid since we are interested in the “paradoxical harm” felt by a candidate (and perhaps their supporters) in an election exhibiting the paradox: if there exists a partition of the electorate such that one of the losing candidates can point to each sub-electorate and claim victory, that candidate can reasonably claim some sort of harm when they ultimately lose the overall election.
A second way our work differs from previous studies is that we use a much larger set of theoretical models as well as empirical data to conduct our analysis. Neither of the previous papers uses any real-world ranked-choice ballot data. [^4] The paper by Courtin et al. does have a section where they apply their results to parliamentary (or House) elections in France, Germany, and the United States, but the authors do not have access to candidate preference rankings. Our use of multiple spatial models, multiple Dirichlet models, and a large dataset of real-world elections (combined with nonparametric bootstrapping analysis) greatly expands previous work.
Thirdly, ours is the first work to provide any analysis of partial ballots in the context of reinforcement paradoxes. Previous work was not concerned with partial ballots, presumably because such work did not analyze any real-world IRV elections.
Finally, both previous papers report extremely small probabilities for observing a reinforcement paradox under IRV. Plassmann and Tideman give a probability of approximately 0.02% when using an electorate of at least 1,000 voters. Courtin et al. report similarly small probabilities. Using the term “Young’s reinforcement axiom” (YRA) synonymously with Young’s notion of consistency (Definition 1), Courtin et al. state, “Our results show that … cases of violations of YRA are rather rare… [[A]]lthough all these [[iterative scoring]] rules violate YRA, the frequencies of violation are very rare. Then, since the violation of this axiom is not so frequent, one should not be unduly worried about its theoretical possibility.” By contrast, we report extremely large reinforcement paradox probabilities, which are generally much larger than the probabilities of observing other classical paradoxes. The reason for this discrepancy is our difference in approach, as outlined above. We would obtain similarly small probabilities if, for example, we were to look at each real election and search for a reinforcement paradox by choosing a partition of the ballots at random. Our probabilities are higher because we consider all possible partitions simultaneously.
7. Conclusion
Both theoretical and empirical analysis suggests that 3-candidate IRV elections are highly susceptible to the reinforcement paradox. As noted above, this is partly because we consider all possible partitions. If we were require the two sub-elections to be similarly-sized, for instance, as they might be in two precincts, there will most likely be fewer instances of the paradox; the same is true if we restrict the sub-elections so that one sub-election is a small perturbation of the original election. Future work could examine how the probability of a reinforcement paradox changes under restrictions of this kind, or to instances involving a larger number of sub-elections.
Future work could also focus on geography-based decompositions of the electorate. Informally, study of the reinforcement paradox is often motivated in geographic terms using a story like this: Suppose a mayoral candidate wins when using only the ballots from the eastern half of a city and wins when using only the ballots from the western half. Then the candidate should win the overall election; a loss would be additively paradoxical. Some of our real-world data comes equipped with precinct information, allowing for geographical analysis. In practice, the number of precincts is often too large to check all possible decompositions of the electorate two precinct subsets, but perhaps heuristic algorithms could be developed to investigate if there are decompositions of the electorate into two sets of precincts, and if these sets of precincts are geographically clustered. The only such example of which we are aware is 2021 Ward 2 City Council IRV election in Minneapolis, in which candidate Yusra Arab wins when using only the ballots from precincts 1 and 5, and wins when using the ballots from all other precincts, but loses the overall election [35]. The number of precincts in Minneapolis’ second ward is small and this example was found using brute force.
Finally, we focus on IRV in this article because of its use in real-world political elections, but it would be interesting to repeat our analysis for other voting methods which are not positional scoring rules. For example, many researchers in the social choice space feel that Condorcet consistent methods are superior to IRV (see Chapter 2 of [13], for example), and thus we would welcome an examination of the susceptibility of Condorcet consistent methods to the reinforcement paradox.